### v1

# Strong Watanabe–Yoshida conjecture for complete intersections

Abstract.
In this paper, we prove the strong form of the Watanabe–Yoshida conjecture for complete intersection singularities in every positive characteristic. In characteristics 2 and 3, we explicitly compute the Hilbert–Kunz functions of the $A_{1}$ and $A_{2}$ singularities.
Introduction
Let $(R,\mathfrak{m})$ be a local Noetherian ring of dimension $d\geq 0$ and characteristic $p>0$. We denote by $\mathfrak{m}^{[p^{e}]}=(a^{p^{e}}:a\in\mathfrak{m})$. The function introduced in [Kun69],
$\text{HK}_{e}(R):=\operatorname{length}_{R}{\left(\frac{R}{\mathfrak{m}^{[p^{e}]}}\right)}$
is called the Hilbert-Kunz function of $R$. Kunz also showed that this function detects singularities of local rings. Monsky [Mon83] proved that the limit
$e_{\text{HK}}(R)=\lim_{e\rightarrow\infty}\frac{\text{HK}_{e}(R)}{p^{ed}}$
always exists, and is greater or equal than one. This limit is called the Hilbert–Kunz multiplicity of $R$. Watanabe and Yoshida [WY05] proved that an unmixed local ring with $e_{\text{HK}}\left(R\right)=1$ has to be regular. This number need not be an integer, but it does not approach 1 arbitrarily when dimension and characteristic are fixed – see [BE04] and [AE08]. The Watanabe–Yoshida conjecture states that the infimum is attained by the quadric hypersurfaces, and also provides a characteristic-free lower bound:
Conjecture ([WY05], [Yos09]).
Let $p>0$ prime and $d\geq 2$. Set $R_{p,d}:=\mathbb{F}_{p}[[x_{0},...,x_{d}]]/(Q_{d}),$ where
$Q_{d}=\begin{cases}x_{0}x_{1}+x_{2}x_{3}+\dots+x_{d-1}x_{d}&\text{if $d$ is odd},\\ x_{0}x_{1}+x_{2}x_{3}+\dots+x_{d-2}x_{d-1}+x_{d}^{2}&\text{if $d$ is even}.\end{cases}$
Let $(R,\mathfrak{m},k)$ be an unmixed non-regular local ring of dimension $d$ and characteristic $p$. Then
  1. (1)
    $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})$,
  2. (2)
    if $e_{\text{HK}}\left(R\right)=e_{\text{HK}}\left(R_{p,d}\right)$, then $\hat{R}\cong R_{p,d}\hat{\otimes}_{\mathbb{F}_{p}}k$, i.e. the completion of $R$ is isomorphic to the coordinate ring of a quadric hypersurface.
  3. (3)
    $e_{\text{HK}}(R_{p,d})\geq 1+c_{d}$,
where $c_{d}$ are the coefficients of the Taylor expansion of $\sec x+\tan x$.
Note that by unmixed we mean that $\dim\hat{R}/\mathfrak{p}=\dim\hat{R}$ for every $\mathfrak{p}\in\text{Ass}(\hat{R})$.
The first two parts are what we refer to as the strong Watanabe–Yoshida conjecture since they not only provide a lower bound for the Hilbert–Kunz multiplicity but also characterize the rings with minimal Hilbert–Kunz multiplicity among complete intersections.
The formula in the third part is often called the weaker inequality, or the weaker part of the conjecture. The bound is motivated by the result of Gessel and Monsky [GM10] that $\lim_{p\rightarrow\infty}e_{\text{HK}}\left(R_{p,d}\right)=1+c_{d}$. Moreover, Yoshida conjectured that $e_{\text{HK}}\left(R_{p,d}\right)$ is a decreasing sequence in $p$. Trivedi showed this for $p$ sufficiently larger than the dimension; see [Tri23].
Different cases of the Watanabe–Yoshida conjecture have been proven, but we still lack a general answer. The following are the existing results in low dimension:
Theorem ([WY05, WY01, AE12, CSA24, Yos09]).
Using the same notation as in the Conjecture above, with $p>2$, and $e(R)$ the (usual) multiplicity of $R$, the following hold.
  1. (1)
    If $d\leq 7$, then $e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(R_{p,d}\right)\geq 1+c_{d}$.
  2. (2)
    If $2\leq d\leq 4$, then $e_{\text{HK}}\left(R\right)=e_{\text{HK}}\left(R_{p,d}\right)$ if and only if $\hat{R}\cong R_{p,d}$.
  3. (3)
    If $e(R)\leq 5$, then $e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(R_{p,d}\right)$.
In 2005, Enescu and Shimomoto proved the first inequality of the conjecture for complete intersections in characteristic $p>2$, and provided a stronger result in characteristic $p>3$:
Theorem ([ES05]).
Let $(R,\mathfrak{m})$ be a non-regular complete intersection of characteristic $p>2$ and dimension $d\geq 1$. Then, $e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(R_{p,d}\right).$ Moreover, if $p>3$, and $\hat{R}\not\cong R_{p,d}$, then
$e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)\geq e_{\text{HK}}\left(R_{p,d}\right),$ (1)
where $S_{p,d}=\overline{\mathbb{F}_{p}}[[x_{0},\dots,x_{d}]]/(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3})$.
However, to our knowledge, it has not been proven that $R_{p,d}$ and $S_{p,d}$ have different Hilbert–Kunz multiplicities. In other words, the Hilbert–Kunz multiplicity distinguishes the $A_{1}$ and $A_{2}$ singularities; see [GK90]. This would directly imply the first two parts of the Watanabe–Yoshida conjecture for complete intersections. Proving this and extending the result to characteristics 2 and 3 is the main result of this paper:
Theorem A (Theorems 2.17 and 4.8).
Let $(R,\mathfrak{m})$ be a non-regular complete intersection of characteristic $p>0$. If $\hat{R}\not\cong R_{p,d}$, then
$e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right),$
where we define $S_{p,d}\coloneqq\mathbb{F}_{p}[[x_{0},...,x_{d}]]/(P_{d})$ with
$P_{d}=\begin{cases}x_{0}x_{1}+\dots+x_{d-1}^{2}+x_{d}^{3}&\text{if $d$ is odd},\\ x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{3}&\text{if $d$ is even},\end{cases}$
for every $p>0$ and $d\geq 2$. In particular, the strong Watanabe–Yoshida conjecture holds for complete intersections.
The cases of characteristic $p>3$ and $2\leq p\leq 3$ are approached slightly differently since in the case of $p=2,3$ we can compute the Hilbert–Kunz function and multiplicity of the $A_{1}$ and $A_{2}$ singularities:
Theorem B (Theorems 3.10, 3.14, 3.15).
For $d\geq 1$,
$e_{\text{HK}}\left(R_{2,d}\right)=\begin{dcases}\frac{2^{m}}{2^{m}-1}&(d=2m-1)\\ \frac{2^{m}+1}{2^{m}}&(d=2m)\\ \end{dcases}\quad\text{and}\quad e_{\text{HK}}\left(S_{2,d}\right)=\begin{dcases}\frac{2^{m-1}+1}{2^{m-1}}&(d=2m-1)\\ \frac{2^{m+1}+1}{2^{m+1}-1}&(d=2m)\\ \end{dcases}$
and
$e_{\text{HK}}\left(R_{3,d}\right)=1+\left(\frac{2}{3}\right)^{d-1}\quad\text{and}\quad e_{\text{HK}}\left(S_{3,d}\right)=1+\frac{3\cdot 2^{d}}{3^{d+1}-2^{d}+(-1)^{d}}.$
Theorem 3.10 confirms Yoshida’s conjecture on the formula of the Hilbert–Kunz function of $e_{\text{HK}}\left(R_{2,d}\right)$. Moreover, with Theorem 3.15, it disproves Conjecture 2.7 in [JNS+23].
We obtain the following as a corollary of the main results of this work.
Theorem C (Theorem 5.1, 5.6).
The strong Watanabe–Yoshida conjecture holds in dimension $d\leq 6$ or multiplicity $e(R)\leq 5$, in characteristic $p>0$. Also if $d=7$ and $p>2$.
Let us now lay out the structure of this paper. Section 1 provides some preliminaries on the Han–Monsky ring since this is the main technique used in the following sections. See [HM93] for details.
Sections 2 and 3 are devoted to showing that the $A_{1}$ and $A_{2}$ singularities indeed have different Hilbert–Kunz multiplicity. We will tackle the characteristic $p>3$ case first in Section 2, and we will solve the characteristics $2$ and $3$ cases in Section 3. The Hilbert–Kunz functions of the $A_{1}$ and $A_{2}$ singularities in characteristic $2$ and $3$ are explicitly computed in that section.
In Section 4 of this paper, we will extend Enescu and Shimomoto’s result to characteristics 2 and 3, in Theorems 4.5 and 4.7, which implies that the strong Watanabe–Yoshida conjecture also holds in those two cases.
Finally, in Section 5, we will revisit the low-dimensional and low-multiplicity cases of the conjecture that have been solved. We will refine them when necessary to obtain the strong Watanabe–Yoshida conjecture, including the characteristic 2 case, from the main results of this work.
1. The Han–Monsky ring.
This section briefly accounts for the Han–Monsky ring based on Yoshida’s unpublished notes [Yos09]. The Han–Monsky ring is the Grothendieck group of finitely generated $k[T]$-modules where the action of $T$ is nilpotent. It was introduced in [HM93] to compute the Hilbert–Kunz function of diagonal hypersurfaces.
Definition 1.1 ($k$-object).
Let $k$ be a field. A $k$-object is a finitely generated $k[T]$-module $M$ such that the action of $T$ over $M$ is nilpotent.
Notation.
We often obtain $k$-objects from taking a zero-dimensional $k$-algebra $R$ and defining the action of $T$ by multiplication by an element $f\in R$. We usually denote the $k$-object as a pair $(R,f)$.
Proposition 1.2 (Definition 1.3, [HM93]).
Let $M,N$ be two $k$-objects. The $k$-vector spaces $M\otimes_{k}N$ and $M\oplus N$ can be regarded as $k$-objects if we define
$T\cdot(m\oplus n)$ $:=Tm\oplus Tn$
$T\cdot(m\otimes_{k}n)$ $:=Tm\otimes_{k}n+m\otimes_{k}Tn.$
Example 1.3.
Take the $k$-object $k[x,y]/(x^{a},y^{b})$ with $T$-action $x^{c}+y^{d}$. Then,
$\frac{k[x,y]}{(x^{a},y^{b})}=\frac{k[x]}{(x^{a})}\otimes\frac{k[y]}{(y^{b})},$
where the $k$-objects on the right side have $T$-action $x^{c}$ and $y^{d}$, respectively. Generally, any $T$-action given by a diagonal equation will split this way.
We introduce the following equivalence relation over the pairs of $k$-objects: Let $(M_{1},N_{1})$ and $(M_{2},N_{2})$ be two pairs of $k$-objects. Then,
[…]
The formal difference $M-N$ denotes the equivalence class containing $(M,N)$. In other words, this is the structure of the Grothendieck ring on the free monoid generated by the equivalence classes of $k$-objects.
Definition 1.4 (Han–Monsky ring).
Let
[…]
and define $+$ and $\cdot$ in $\Gamma_{k}$ as follows:
$(M_{1}-N_{1})+(M_{2}-N_{2})$ $=(M_{1}\oplus M_{2})-(N_{1}\oplus N_{2}),$
$(M_{1}-N_{1})\cdot(M_{2}-N_{2})$ $=\left[(M_{1}\otimes M_{2})\oplus(N_{1}\otimes N_{2})\right]-\left[(M_{1}\otimes N_{2})\oplus(M_{2}\otimes N_{1})\right].$
This way we have a commutative ring with unity $k-0$, called the Han–Monsky ring over $k$.
Usually, there is no ambiguity in denoting $\Gamma=\Gamma_{k}$, because $\Gamma$ only depends on the characteristic of the field, which is usually clear from the context.
Definition 1.5.
Let $\delta_{i}:=k[T]/(T^{i}),$ for every $i\geq 1$. We call these elements the $\delta$-basis.
The fact that the $\delta$-basis indeed forms a $\mathbb{Z}$-basis is a rewording of the structure theorem for finitely generated modules over a PID:
Corollary 1.6.
The abelian group $\Gamma$ is a free $\mathbb{Z}$-module where $\delta_{i}$ is a free $\mathbb{Z}$-basis.
Remark.
Let $k[x]/(x^{a})$ be a $k$-object with $T$-action $x^{c}$. Let $a=ck+r$. Then, one can decompose $k[x]/(x^{a})$ as $(c-r)\delta_{k}+r\delta_{k+1}$. Looking back at Example 1.3, if $k[y]/y^{b}$ has $T$-action $x^{d}$ and $b=dl+s$, then $k[x,y]/(x^{a},y^{b})$ with $T$-action $x^{c}+y^{d}$ decomposes in the Han–Monsky ring as $\left((b-r)\delta_{k}+r\delta_{k+1}\right)\left((d-s)\delta_{l}+s\delta_{l+1}\right)$. Therefore, understanding the products $\delta_{i}\delta_{j}$ is crucial to get to a decomposition over the $\delta$-basis.
To understand the multiplication table of $\Gamma$, a different $\mathbb{Z}$-basis is introduced, which has a simpler behavior.
Definition 1.7 ($\lambda$-basis).
Let $\lambda_{i}:=(-1)^{i}(\delta_{i+1}-\delta_{i}),i\geq 0$, noting that $\delta_{0}=0$ in $\Gamma$.
Remark.
Note that $\delta_{i}=\lambda_{0}-\lambda_{1}+\dots+(-1)^{i}\lambda_{i-1},\text{ for }i\geq 1.$
To finish, we define a $\mathbb{Z}$-linear function over $\Gamma$.
Definition 1.8.
Let $M-0=\bigoplus\delta_{i}^{\oplus a_{i}}$, and define
$\alpha(M-0):=\dim_{k}M/TM=\sum_{i=1}^{\infty}a_{i}$
Extending by linearity to all pairs $M-N$, we obtain a $\mathbb{Z}$-linear function $\alpha:\Gamma\rightarrow\mathbb{Z}$. For $s,a_{i},b_{i}\in\mathbb{Z}$ with $s\geq 1$, $a_{i}\geq 1$, and $b_{i}\geq 0$, we denote
$D_{k}(a_{1},\dots,a_{s}):=\alpha(\prod_{i=1}^{s}\delta_{a_{i}})\hskip 40.00006pt\ell_{k}(b_{1},\dots,b_{s}):=\alpha(\prod_{i=1}^{s}\lambda_{a_{i}}).$
Remark.
Note that $\alpha(\delta_{j})=1,\forall j\geq 1$, and
$\alpha(\lambda_{i})=\begin{cases}1&(i=0),\\ 0&(i\geq 1).\end{cases}$
Now, let $R=k[[x_{1},\dots,x_{s}]]/(f)$ be a characteristic $p>0$ hypersurface. Then $\text{HK}_{e}(R)=\dim_{k}\left(k[[x_{1},\dots,x_{s}]]/(f)+\mathfrak{m}^{[p^{e}]}\right).$ Observe that we can regard $\mu_{e}:=k[[x_{1},\dots,x_{s}]]/\mathfrak{m}^{[p^{e}]}$ as a $k$-object with $T$-action $f$, so that $\text{HK}_{e}(R)=\dim_{k}(\mu_{e}/T\mu_{e}).$ If $\mu_{e}=\sum a_{i}\delta_{i}$ with the given $k$-object structure, then $\text{HK}_{e}(R)=\sum a_{i}$.
1.1. Arithmetic of the Han–Monsky ring.
The Han–Monsky ring $\Gamma$ can be filtered by subrings $\Lambda_{e}\subset\Lambda_{e+1}$:
Theorem-Definition 1.9 (Theorem 3.2, [HM93]).
The additive subgroup of $\Gamma$ generated by the $\lambda_{i}$ for $10\leq i\leq p^{e}-$ 1 is closed under multiplication. We call this ring $\Lambda_{e}$.
Theorem 1.10 (Theorem 2.5, [HM93]).
Let $i\leq j\leq p-1$. Then $\lambda_{i}\lambda_{j}=\sum_{j-i}^{\min{i+j,2p-2-i-j}}\lambda_{k}$.
$\cdot$ $\lambda_{0}$ $\lambda_{1}$
$\lambda_{0}$ $\lambda_{0}$ $\lambda_{1}$
$\lambda_{1}$cdot $\lambda_{0}$
$\cdot$ $\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$
$\lambda_{0}$ $\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$
$\lambda_{1}$cdot $\lambda_{0}+\lambda_{1}+\lambda_{2}$ $\lambda_{1}$
$\lambda_{2}$cdot $\cdot$ $\lambda_{0}$
Table 1. Multiplication table for the $\lambda$-basis of $\Lambda_{1}$ in characteristic $2$ and $3$, respectively.
Proposition 1.11.
$\delta_{p}^{2}=p\delta_{p}$.
Proof.
First part of Lemma 3.3 in [HM93]. ∎
Lemma 1.12 (Lemma 3.4, [HM93]).
Let $q=p^{n}$, $n\geq 1$. Take $10\leq i\leq q-$ 1 . Then, in $\Gamma$,
  1. (a)
    $\lambda_{i}\lambda_{qj}=\lambda_{qj+i}$;
  2. (b)
    $\lambda_{i}\lambda_{qj-1}=\lambda_{qj-i-1}$.
Corollary 1.13.
Let $0\leq a<p^{n+1},$ and $a=a_{0}+a_{1}p+a_{2}p^{2}+\dots+a_{n}p^{n}$ be its expansion in base $p$, $0\leq a_{i}<p$. Then, we have the following decomposition:
$\lambda_{a}=\prod_{k\geq 0}^{n}\lambda_{a_{k}p^{k}}.$
Proof.
We proceed by induction on $n$. For $n=0$, there is nothing to do. Now, observe that
$a_{0}+pa_{1}+\dots+p^{n}a_{n}=(a_{0}+pa_{1}+\dots+p^{n-1}a_{n-1})+p^{n}a_{n},$
where $a_{0}+pa_{1}+\dots+p^{n-1}a_{n-1}<p^{n}.$ Thus, the lemma above gives
$\lambda_{a_{0}+pa_{1}+\dots+p^{n-1}a_{n-1}+p^{n}a_{n}}=\lambda_{a_{0}+pa_{1}+\dots+p^{n-1}a_{n-1}}\lambda_{p^{n}a_{n}}.$
This result implies that to understand the multiplication table of $\Gamma$, it is enough to understand the multiplication table of the $\{\lambda_{qi}\}_{0\leq i\leq q-1}$ for each $q=p^{e}$. For characteristic 2, the following result solves it:
Lemma 1.14 (Theorem 3.7, [HM93]).
Given $q=2^{n},$ then $\lambda_{q}^{2}=\lambda_{0}$.
We finish the section with the Han–Monsky ring with rational coefficients:
Definition 1.15 (Definition 4.8, [HM93]).
Let $a\geq 0$ be a real number. In $\Gamma^{\prime}:=\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$, set
$\delta_{a}:=(1-z)\delta_{r}+z\delta_{r+1},\text{ where }a=r+z,r\in\mathbb{Z}_{\geq 0},z\in[0,1).$
We extend $\alpha$ to $\Gamma^{\prime}$ by $\mathbb{R}$-linearity, and denote $D(a_{i},\dots,a_{i}):=\alpha(\prod\delta_{a_{i}})$. Note that $\delta_{0}=0$.
2. Strong Watanabe–Yoshida for Complete Intersections of characteristic $p>3$
For the rest of the section, let $p>3$ and $d\geq 2$, unless otherwise stated. We will deal with the cases when $p=2$ and $p=3$ in the next section.
Definition 2.1 ($A_{1}$ and $A_{2}$ singularities, [GK90]).
Let $k$ be a field of positive characteristic $p$. The rings $R_{p,d}:=k[[x_{0},...,x_{d}]]/(Q_{d})$ where
$Q_{d}:=\begin{cases}x_{0}x_{1}+\dots+x_{d-1}x_{d},&\text{if $d$ is odd},\\ x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{2},&\text{if $d$ is even},\end{cases}$
are the $A_{1}$ singularities. The rings $S_{p,d}\coloneqq k[[x_{0},...,x_{d}]]/(P_{d})$ where
$P_{d}=\begin{cases}x_{0}x_{1}+\dots+x_{d-1}^{2}+x_{d}^{3}&\text{if $d$ is odd},\\ x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{3}&\text{if $d$ is even}.\end{cases}$
are the $A_{2}$ singularities.
Note that for $p\geq 3$,
$R_{p,d}\cong\frac{k[[x_{0},\dots,x_{d}]]}{(x_{0}^{2}+\dots+x_{d}^{2})},\ S_{p,d}\cong\frac{k[[x_{0},\dots,x_{d}]]}{(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3})},$
so the $A_{1}$ and $A_{2}$ singularities are diagonal hypersurfaces in characteristic $p\geq 3$. This is crucial to prove the main result of this section, the inequality $e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right)$, since we will use Han and Monsky’s results in [HM93], the so-called Han–Monsky algorithm.
2.1. The Hilbert–Kunz function of a diagonal hypersurface
Let us recap some results in Section 5 of [HM93]. Let $R$ denote the ring of power series at the origin of a diagonal hypersurface:
$R:=\frac{k[[x_{0},\dots,x_{s}]]}{(x_{0}^{d_{0}}+\dots+x_{s}^{d_{s}})}.$
Lemma 2.2 (Lemma 5.6, [HM93]).
Let $q=p^{e}$. The $k$-object $k[[x_{0},\dots,x_{s}]]/(x_{0}^{q},\dots,x_{s}^{q})$ with $T$-action $x_{0}^{d_{0}}+\dots+x_{s}^{d_{s}}$ is $\prod_{i=0}^{s}\delta_{q/d_{i}}$ in $\Gamma^{\prime}$. Therefore,
$\text{HK}_{e}(R)=d_{0}\dots d_{s}D\left(\frac{q}{d_{0}},\dots,\frac{q}{d_{s}}\right).$
This result expresses that we can understand the Hilbert–Kunz function of diagonal hypersurfaces by studying the function $D(a_{0},\dots,a_{s})$. The Han–Monsky algorithm is the systematic method for computing a formula for $e\mapsto D(q/d_{i},\dots,q/d_{s})$ for $e$ large enough.
This algorithm starts with three data: the characteristic $p$, the vector of inverses of the exponents of the diagonal hypersurface $\beta=(1/d_{0},\dots,1/d_{s})$, and a period $\mu$ of our choice. Should the period be chosen wisely – which can always be done – we will be under the hypotheses of Theorem 2.4 that spells out the Hilbert–Kunz function of the given diagonal hypersurface for every large $e$ in terms of these two data. To work with said theorem, we need to produce a series of vectors out of $p$, $\mu$, and $\beta$:
Definition 2.3.
Let $\beta$ be a vector of positive real numbers and $p$ be a prime. Let $\mu$ be a positive integer that we call the period. We obtain the following data associated to these $\beta,\mu$, and $p$.
  1. (1)
    Let $r\in\mathbb{Z}_{\geq 0}^{s}$ and $z\in[0,1)^{s+1}$ such that $\beta=r+z$.
  2. (2)
    Let $R\in\mathbb{Z}_{\geq 0}^{s}$ and $z^{\prime}\in[0,1)^{s+1}$ such that $p^{\mu}\beta=R+z^{\prime}$.
  3. (3)
    Let $t:=p^{\mu}z-z^{\prime}=t\in\mathbb{Z}_{\geq 0}^{s}$.
We denote the coordinates as $r=(r_{0},\dots,r_{s})$.
Let $v$ and $w$ be arbitrary vectors. We say that $v$ is equivalent to $w$, denoted $v\sim w$, if we can obtain it from $w$ by replacing $v_{i}$ by $1-v_{i}$ for an even number of indices $i$ (so in particular, if they are equal to each other). We also introduce the notation $v^{*}:=(1-v_{0},v_{1},\dots,v_{s})$.
Finally, we define the function $\ell^{\sharp}$:
=$\ell^{\sharp}=\begin{cases}\ell_{k}(t)&\text{if }\sum r_{i}\text{even,}\\ \ell_{k}(p^{\mu}-1-t_{0},t_{1},\dots,t_{s})&\text{if }\sum r_{i}\text{ odd.}\end{cases}$
We will refer to the elements $r,z,R,z^{\prime},t$ and $\ell^{\sharp}$ as the data associated to $\beta,\mu$ and $p$.
Theorem 2.4 (Theorem 5.3 and remark, [HM93]).
Let $\beta$ be a vector of positive real numbers, and $\mu$ a positive integer, and let $r,z,R,z^{\prime}$, and $t$ be the data associated to $\beta,\mu$, and $p$.
Suppose that either one of these two sets of conditions is satisfied by the data:
  1. (1)
    $z^{\prime}\sim z$ and $\sum R_{i}\equiv\sum r_{i}\mod{2}$, or
  2. (2)
    $z^{\prime}\sim z^{*}$ and $\sum R_{i}\not\equiv\sum r_{i}\mod{2}$.
Then, there is a rational number $C$ and rational numbers $\Delta_{0},\Delta_{1},\Delta_{2},\dots$ such that
$D(p^{e}\beta)=Cp^{de}+\Delta_{e},$
and $\Delta_{e+\mu}=\ell^{\sharp}\Delta_{e}$ for all $e\geq 0$, where $\ell^{\sharp}$ was defined in 2.3.
We note that given any $\beta$ vector with non-negative rational numbers, there always exists a period such that the first set of conditions in the hypotheses of Theorem 2.4 is satisfied, as it is shown in the proof of Theorem 5.3 of [HM93] – Han and Monsky provide the second set of conditions as a mere remark. However, to prove the main result of the section, we will exploit both sets of conditions.
2.2. The Hilbert–Kunz multiplicity of $A_{1}$ and $A_{2}$ singularities, $p>3$.
It turns out that in the case that the diagonal hypersurface is an $A_{1}$ or an $A_{2}$ singularity, we can choose the period to be 1 for every $p>3$ and $d\geq 2$:
Lemma 2.5.
Let $\beta=(\frac{1}{2},\dots,\frac{1}{2},\frac{1}{3})$, $\mu=1$, and $p>3$. Then, the data $r,z,R,z^{\prime},t,\ell^{\sharp}$ associated to this pair satisfy the hypotheses of Theorem 2.4. Also, $\sum r_{i}=0$, so $\ell^{\sharp}=\ell_{k}(t)$. The same follows taking $\beta=(\frac{1}{2},\dots,\frac{1}{2})$, $\mu=1$ and $p>2$.
Proof.
Let us compute $r,z$ and $R$ first, according to Definition 2.3:
$r=(0,\dots,0),z=\beta,R=(\left\lfloor\frac{p}{2}\right\rfloor,\dots,\left\lfloor\frac{p}{2}\right\rfloor,\left\lfloor\frac{p}{3}\right\rfloor).$
Firstly, observe that $z=z^{*}$, given that $z_{0}=\frac{1}{2}$. Thus, the hypotheses of the Theorem will be satisfied regardless of the parity of $\sum R_{i}$ once we prove that $z^{\prime}$ is equivalent to $z$.
When computing $z^{\prime}$, two cases arise: either $p\equiv 1\mod{3}$ or $p\equiv 2\mod{3}$.
  1. (1)
    If $p\equiv 1\mod{3}$, then $z^{\prime}=(\frac{1}{2},\dots,\frac{1}{2},\frac{1}{3})=z$.
  2. (2)
    If $p\equiv 2\mod{3}$, then $z^{\prime}=(\frac{1}{2},\dots,\frac{1}{2},\frac{2}{3})$. In this case, given that $s\geq 1$, and the first $s$ entries of $\beta$ are $1/2$, one sees that $z^{\prime}$ is equivalent to $z$ by observing that $z^{\prime}=(1-z_{0},z_{1},\dots,z_{s-1},1-z_{s})$.
The result follows.
As for $\beta=(1/2,\dots,1/2)$, $\mu=1$ and $p>2$, it also happens that $z=z^{*}$, and so we do not have to bother about the parity of $\sum R_{i}$ either. Moreover, $p\equiv 1\pmod{2}$ for every $p>2$, so $z=z^{\prime}$. ∎
The fact that the period can be chosen to be $1$ is useful because it will allow us to exploit the simplicity of the multiplication table of $\Lambda_{1}$. It will be for $S_{p,d}$ that we will need to use the second set of conditions in 2.4. These observations are the content of the next lemma.
Theorem 2.6.
For all $p>2$ and $d\geq 0$,
$e_{\text{HK}}(R_{p,d})=1+\frac{\text{HK}_{1}(R_{p,d})-p^{d}}{p^{d}-\ell^{\sharp}(R_{p,d})},$
where $\ell^{\sharp}(R_{p,d})=\alpha(\lambda_{a}^{d+1})$ for $a=\lfloor\frac{p}{2}\rfloor$. Similarly, for all $p>3$ and $d\geq 0$, it follows that
$e_{\text{HK}}(S_{p,d})=1+\frac{\text{HK}_{1}(S_{p,d})-p^{d}}{p^{d}-\ell^{\sharp}(S_{p,d})},$
where $\ell^{\sharp}(S_{p,d})=\alpha(\lambda_{a}^{d}\lambda_{b})$ for $a=\lfloor\frac{p}{2}\rfloor$ and $b=\lfloor\frac{p}{3}\rfloor$.
Proof.
By Lemma 2.2, we know that $\text{HK}_{e}(S_{p,d})=2^{d}3D(p^{e}\beta)$. Then, Theorem 2.4 yields that there exist rational numbers $C$, $\ell^{\sharp}_{S}$, and $\Delta_{e}$ such that
$2^{d}3D(p^{e}\beta)=2^{d}3Cp^{de}+2^{d}3\Delta_{e},$
where $\Delta_{e+1}=\ell^{\sharp}_{S}\Delta_{e}$ for all $e\geq 0$. Note that $e_{\text{HK}}\left(S_{p,d}\right)=2^{d}3C$ and let $c_{1}=-2^{d}3\Delta_{0}$, so
$\text{HK}_{e}(S_{p,d})=e_{\text{HK}}\left(S_{p,d}\right)p^{de}-c_{1}(\ell^{\sharp}_{S})^{e},e\geq 0.$
The same argument applies to $R_{p,d}$, for which we obtain
$\text{HK}_{e}(R_{p,d})=e_{\text{HK}}\left(R_{p,d}\right)p^{de}-c_{1}(\ell^{\sharp}_{R})^{e},e\geq 0.$
Since $\text{HK}_{0}(R)=1$ for every ring, the computations above imply that
$\begin{dcases}e_{\text{HK}}\left(S_{p,d}\right)-c_{1}=1,\\ e_{\text{HK}}\left(S_{p,d}\right)p^{d}-c_{1}\ell^{\sharp}=\text{HK}_{1}(S_{p,d}).\end{dcases}$
Replacing $e_{\text{HK}}\left(S_{p,d}\right)=1+c_{1}$ on the second equation, we obtain
$p^{d}+c_{1}p^{d}-c_{1}\ell^{\sharp}=\text{HK}_{1}(S_{p,d})\Rightarrow c_{1}=\frac{\text{HK}_{1}(S_{p,d})-p^{d}}{p^{d}-\ell^{\sharp}}.$
The same argument applies to $e_{\text{HK}}\left(R_{p,d}\right)$. ∎
Remark.
The $A_{1}$ and $A_{2}$ singularities are not the only ones for which the formula above works. We can easily see that we get the same result for other diagonal hypersurfaces, for example, whenever $d_{0}=2$ and $p\equiv\pm 1\mod{d_{i}}$ for every $i$. In this case, it happens that $z=z^{*}$, and therefore $z^{\prime}$ is equivalent to both $z$ and $z^{*}$, and both sets of conditions in Theorem 2.4 are satisfied. This is the case, for instance, of $A_{3}$ singularities with $p\geq 5$. The Hilbert–Kunz multiplicity of an $A_{n}$ singularity (see [GK90]) verifies the formula in the previous theorem whenever $p\equiv\pm 1\pmod{n+1}$.
However, the formula does not work in full generality – take the $A_{4}$ surface singularity in characteristic 7, which has the equation $x^{2}+y^{2}+z^{5}$. The formula would give the value $43/25$, but it has Hilbert–Kunz multiplicity $2-1/5=9/5$ (see, for instance, Example 4.1 [WY01]).
2.3. The powers of $\gamma$ in the Han–Monsky ring.
In this section, we prove some necessary technical corollaries in the following section and involve computations in the Han–Monsky ring.
Notation.
Let $p>2$ be a prime and $k$ a field of characteristic $p$. For the rest of the section, we will denote the $k$-object $\gamma:=\left(\frac{k[x]}{(x^{p})},x^{2}\right)$ and for $p>3$ prime, we will write $\eta:=\left(\frac{k[x]}{(x^{p})},x^{3}\right)$. Observe that $\gamma\in\Lambda_{1}$. We will use the following notation for the coordinates of the powers of $\gamma$ over the $\lambda$-basis:
$\gamma^{d}=a_{0}^{(d)}\lambda_{0}+\dots+a_{p-1}^{(d)}\lambda_{p-1}.$
Proposition 2.7.
Let $k$ be a field of characteristic $p\geq 3$. Set $a=(p-1)/2$. Then,
$\gamma=\delta_{a}+\delta_{a+1}=2\lambda_{0}-2\lambda_{1}+\dots+(-1)^{a-1}2\lambda_{a-1}+(-1)^{a}\lambda_{a}.$
The point of this section is to prove that the first $a+1$ coefficients of $\gamma^{d}$ form a monotone sequence – proving this is the goal of the following propositions.
Lemma 2.8 ([Yos09]).
Let $k\geq a$. Then,
$\gamma\lambda_{k}=\begin{dcases}(-1)^{k}\left(2\sum_{i=0}^{a-k-1}(-1)^{i}\lambda_{i}+\sum_{i=a-k}^{a+k}(-1)^{i}\lambda_{i}\right)&(k<a);\\ (-1)^{a}\delta_{p}&(k=a),\end{dcases}$
Proof.
Explicit computation using Lemma 1.10 and Proposition 2.7. This formula is provided in matrix form in Section 4 of Yoshida’s notes [Yos09]. ∎
Remark.
Observe that for $p-1\geq k\geq a+1$, $\lambda_{k}\gamma=\lambda_{p-1}\lambda_{p-1-k}\gamma$, by 1.12(b).
Lemma 2.9.
Let $p>3$ and $d\geq 1$. Let $a,b,c\in\mathbb{Z}_{\geq 0}$ such that $p=3b+c$ and $p=2a+1$, and $c<3$. Then,
$\alpha(\gamma^{d}\eta)$ $=3a_{0}^{(d)}+\dots+3a_{b-1}^{(d)}+ca_{b}^{(d)},$
$\alpha(\gamma^{d+1})$ $=2a_{0}^{(d)}+\dots+2a_{a-1}^{(d)}+a_{a}^{(d)}.$
Proof.
Observe that, by definition and Lemma 2.8,
$\alpha(\gamma^{d+1})=\sum_{i=0}^{p-1}(-1)^{i}a_{i}^{(d)}\alpha(\lambda_{i}\gamma)=\sum_{i=0}^{a-1}(-1)^{i}a_{i}^{(d)}(-1)^{i}2+(-1)^{i}a_{a}^{(d)}(-1)^{i}.$
Lemma 2.10.
For every $d\geq 1$, $\gamma^{d}(\lambda_{p-1}+\lambda_{0})=2\cdot p^{d-1}\delta_{p}.$ Equivalently, if we denote $\gamma^{d}=\sum_{i=0}^{p-1}(-1)^{i}a_{i}^{(d)}\lambda_{i}$, then
$a_{i}^{(d)}+a_{p-i}^{(d)}=2p^{d-1}.$
In particular, this implies that $a_{a}^{(d)}=p^{d-1}$, for $a=\frac{p-1}{2}$.
Proof.
One can check by Lemma 2.8 that $\gamma(\lambda_{0}+\lambda_{p-1})=2\delta_{p}$. Now, $(\lambda_{0}+\lambda_{p-1})^{2}=2(\lambda_{0}+\lambda_{p-1})$ by 1.12. By 1.11, the result follows. ∎
Lemma 2.11 (Theorem 4.1, [Yos09]).
For $d\geq 2$,
$\gamma^{d}-p^{d-1}\delta_{p}=2\sum_{r=0}^{a-1}((-1)^{r}\gamma^{d-1}\lambda_{r}-p^{d-2}\delta_{p}).$
Equivalently, for $k=1,\dots,a$
$a_{a-k}^{(d)}-p^{d-1}=2\sum_{r=a-k}^{a-1}(a^{(d-1)}_{a-r}-p^{d-2}).$
Proof.
Lemma 2.10 applied twice yields that
$\gamma^{d}-p^{d-1}\delta_{p}$ $=p^{d-1}\delta_{p}-\gamma^{d}\lambda_{p-1}=p^{d-1}\delta_{p}-\gamma(2p^{d-2}\delta_{p}-\gamma^{d-1})$
$=p^{d-1}\delta_{p}+2\sum_{r=0}^{a-1}(-1)^{r}\lambda_{r}(\gamma^{d-1}-2p^{d-2}\delta_{p})+(-1)^{a}\lambda_{a}(\gamma^{d-1}-2p^{d-2}\delta_{p}),$
and from Lemma 1.10 one can check that $(-1)^{r}\lambda_{r}\delta_{p}=\delta_{p}$, which yields
$\gamma^{d}-p^{d-1}\delta_{p}$ $=p^{d-1}\delta_{p}+2\sum_{r=0}^{a-1}((-1)^{r}\lambda_{r}\gamma^{d-1}-2p^{d-2}\delta_{p})+((-1)^{a}\lambda_{a}\gamma^{d-1}-2p^{d-2}\delta_{p})$
$=p^{d-1}\delta_{p}-(p-1)p^{d-2}\delta_{p}+2\sum_{r=0}^{a-1}((-1)^{r}\lambda_{r}\gamma^{d-1}-p^{d-2}\delta_{p})+((-1)^{a}\lambda_{a}\gamma^{d-1}-2p^{d-2}\delta_{p}).$
To finish, note that $p^{d-1}\delta_{p}-(p-1)p^{d-2}\delta_{p}=p^{d-2}\delta_{p},$ and adding the term in the rightmost side of the formula above,
$p^{d-2}\delta_{p}+(-1)^{a}\lambda_{a}\gamma^{d-1}-2p^{d-2}\delta_{p}=(-1)^{a}\lambda_{a}\gamma^{d-1}-p^{d-2}\delta_{p}.$
It follows from Lemma 2.10 and the remark above that this is zero, so the result follows. ∎
Remark.
Observe that, for $d\geq 2$, $2\sum_{r=0}^{a-1}((-1)^{r}\gamma^{d-1}\lambda_{r}-p^{d-2}\delta_{p})=2(\gamma^{d-1}-p^{d-2}\delta_{p})\delta_{a}.$ Applying the result recursively, one obtains that $\gamma^{d}-p^{d-1}\delta_{p}=2^{d-1}(\gamma-\delta_{p})\delta_{a}^{d-1}$, for $d\geq 1$, which is an interesting identity on its own.
From this lemma, we can now prove the result we were aiming for. We prove something slightly stronger, which is that the sequence is always strictly increasing for $d\geq 1$, but we will only use the fact that it is increasing.
Proposition 2.12.
Fix $d\geq 1$. Then, the sequence $a_{i}^{(d)}$ is decreasing on $0\leq i\leq a$. It is strictly decreasing for $d\geq 2$.
Proof.
Let $k=0,\dots,a-1$. Then, by Lemma 2.11,
$a_{a-k}^{(d)}-a_{a-k-1}^{(d)}=2\sum_{r=a-k}^{a}(a_{a-r}^{(d-1)}-p^{d-2})-2\sum_{r=a-k-1}^{a}(a_{a-r}^{(d-1)}-p^{d-2})=2(a_{k+1}^{(d-1)}-p^{d-2}).$
Since $p^{d-2}=a_{a}^{(d-1)}$, the result follows from Proposition 2.7 and induction on $d$. ∎
2.4. The inequality.
We finally conclude in this section that the Hilbert–Kunz multiplicity of the $A_{1}$ and $A_{2}$ singularities are different.
Theorem 2.13 (Theorem 4.6, [HM93]).
If $d\geq 2$, $q=p^{e}$ and $k_{i}\leq q-1$, then $\alpha(\prod_{i=0}^{d}\lambda_{k_{i}})\leq q^{d-2}$. In particular, $\ell^{\sharp}(S_{p,d})\leq p^{d-2}$.
Proposition 2.14.
For $d\geq 2$ and $p>3$ prime, $\text{HK}_{1}(S_{p,d})-\text{HK}_{1}(R_{p,d})\geq\text{HK}_{1}(R_{p,d-1})-p^{d-1}$.
Proof.
Let $a,b,c$ like in 2.9. Since $d$ is fixed throughout the proof, there is no ambiguity in denoting $\gamma^{d}=\sum_{i=0}^{p-1}a_{i}\lambda_{i}$. By Lemma 2.9, rearranging the coefficients we obtain:
$\text{HK}_{1}(S_{p,d})=\alpha(\gamma^{d}\eta)=\sum_{k=1}^{p}a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}\quad\text{and}\quad\text{HK}_{1}(R_{p,d})=\alpha(\gamma^{d+1})=\sum_{k=1}^{p}a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}.$
Thus,
$\text{HK}_{1}(S_{p,d})-\text{HK}_{1}(R_{p,d})=\sum_{k=1}^{p}\left(a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\right).$
Now, it happens that $\lfloor\frac{k+1}{2}\rfloor\geq\lfloor\frac{k+2}{3}\rfloor$, and the equality only holds for $k=1,2$ or $4$. This implies that $\lfloor\frac{k+1}{2}\rfloor>\lfloor\frac{k+2}{3}\rfloor$ for $k=3$ and $p\geq k\geq 5$. Hence, by Proposition 2.12
$a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\neq 0$
for $k=3$ and $p\geq k\geq 5$. So,
$\sum_{k=1}^{p}\left(a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\right)=(a_{0}-a_{1})+\sum_{k=5}^{p}\left(a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\right).$
where every term of the sum on the right side is non-zero. Note that
$a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\geq a_{\left\lfloor\frac{k+1}{2}\right\rfloor-2}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}$
therefore
$\sum_{k=5}^{p}\left(a_{\left\lfloor\frac{k+2}{3}\right\rfloor-1}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\right)$ $\geq\sum_{k=5}^{p}\left(a_{\left\lfloor\frac{k+1}{2}\right\rfloor-2}-a_{\left\lfloor\frac{k+1}{2}\right\rfloor-1}\right)$
$=2(a_{1}-a_{2})+2(a_{2}-a_{3})+\dots+2(a_{a-2}-a_{a-1})+(a_{a-1}-a_{a})$
$\geq(a_{1}-a_{2})+(a_{2}-a_{3})+\dots+(a_{a-1}-a_{a})=a_{1}-a_{a}$
Thus, to sum up, we just proved that
$\text{HK}_{1}(S_{p,d})-\text{HK}_{1}(R_{p,d})$ $\geq(a_{0}-a_{1})+(a_{1}-a_{a})=a_{0}-a_{a}=\text{HK}_{1}(R_{p,d-1})-p^{d-1}.$
Proposition 2.15.
For $d\geq 2$ and $p>3$ prime, $p\text{HK}_{1}(R_{p,d-1})\geq\text{HK}_{1}(S_{p,d})$.
Proof.
We use the same notation in 2.14. Then, by 2.12,
$\text{HK}_{1}(S_{p,d})=3a_{0}+\dots+3a_{b-1}+ca_{b}\leq 3a_{0}+\dots+3a_{0}+ca_{0}=pa_{0}=p\text{HK}_{1}(R_{p,d-1}).$
Theorem 2.16.
For $p>3$ prime and $d\geq 2$, $e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right)$.
Proof.
By Theorem 2.6 and Corollary 2.13,
$e_{\text{HK}}\left(S_{p,d}\right)=\frac{\text{HK}_{1}(S_{p,d})-p^{d}}{p^{d}-\ell^{\sharp}(S_{p,d})}\geq\frac{\text{HK}_{1}(S_{p,d})-p^{d}}{p^{d}},$
$e_{\text{HK}}\left(R_{p,d}\right)=\frac{\text{HK}_{1}(R_{p,d})-p^{d}}{p^{d}-\ell^{\sharp}(R_{p,d})}\leq\frac{\text{HK}_{1}(R_{p,d})-p^{d}}{p^{d}-p^{d-2}},$
From Propositions 2.14 and 2.15, it follows that
$p^{2}(\text{HK}_{1}(S_{p,d})-\text{HK}_{1}(R_{p,d}))\geq p^{2}(\text{HK}_{1}(R_{p,d-1})-p^{d-1})>\\ >p(\text{HK}_{1}(R_{p,d-1})-p^{d-1})\geq\text{HK}_{1}(S_{p,d})-p^{d},$ (2)
which implies that
$\frac{\text{HK}_{1}(S_{p,d})-p^{d}}{p^{d}}>\frac{\text{HK}_{1}(R_{p,d})-p^{d}}{p^{d}-p^{d-2}}.$
Theorem 2.17 (Strong Watanabe–Yoshida conjecture for complete intersections, $p>3$).
Let $(R,\mathfrak{m})$ be a non-regular complete intersection of characteristic $p>3$ and dimension $d\geq 2$. If $\hat{R}\not\cong R_{p,d}$, then
$e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right),$
where $S_{p,d}\coloneqq\overline{\mathbb{F}_{p}}[[x_{0},...,x_{d}]]/(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3})$.
Proof.
The second inequality is Theorem 2.16. As for the first, let $K|k$ be an uncountable field extension of $k$. Since field extensions are faithfully flat, then $e_{\text{HK}}\left(R\right)=e_{\text{HK}}\left(R\otimes_{k}K\right)$. Now, by Theorem 4.3 in [ES05], there exists some $F\in K[[x_{0},\dots,x_{d}]]$ such that
$e_{\text{HK}}\left(R\otimes_{k}K\right)\geq e_{\text{HK}}\left(K[[x_{0},\dots,x_{d}]]/(F)\right).$
Finally, by Theorem 3.4 in [ES05], we have that $e_{\text{HK}}\left(K[[x_{0},\dots,x_{d}]]/(F)\right)\geq e_{\text{HK}}\left(S_{p,d}\right).$
3. HK functions of the $A_{1}$ and $A_{2}$ singularities in characteristics 2 and 3.
3.1. $\sigma$-basis
We introduce a basis of idempotent elements of $\Lambda_{e}\otimes_{\mathbb{Z}}\mathbb{Q}$, which we will call the $\sigma$-basis. This will be the main tool to compute the Hilbert–Kunz functions of the $A_{1}$ and $A_{2}$ singularities.
Definition 3.1 ($\sigma$-elements).
Let $e\geq 1$ and $10\leq k\leq 2^{e}-$ 1. Write $k$ in binary as $k=k_{0}+2k_{1}+\dots+2^{e-1}k_{e-1}$, where $k_{i}\in\{0,1\}$. We define the $\sigma$-basis as:
$\sigma_{k,e}:=\frac{1}{2^{e}}\prod_{j=0}^{e-1}(\lambda_{0}+(-1)^{k_{j}}\lambda_{2^{j}})\in\Lambda_{e}\otimes_{\mathbb{Z}}\mathbb{Q}$
where $\lambda_{i}$ are the elements of the $\lambda$-basis.
Proposition 3.2.
Let $e\geq 1$. The $\sigma$-elements have the following properties:
  1. (a)
    (Idempotency and orthogonality) $\sigma_{k,e}^{2}=\sigma_{k,e}$ and $\sigma_{k,e}\sigma_{l,e}=0$ for $k\neq l$.
  2. (b)
    If $\eta=\sum_{k=0}^{2^{e}-1}a_{k}\sigma_{k,e}$, then $\sigma_{k,e}\eta=a_{k}\sigma_{k,e}$.
  3. (c)
    (Interaction between $\sigma$-bases) Let $e\leq e^{\prime}$,
    $\sigma_{l,e^{\prime}}\sigma_{k,e}=\begin{cases}\sigma_{l,e^{\prime}}&\text{ if }l=k\text{ or }|l-k|\geq 2^{e},\\ 0&\text{otherwise.}\end{cases}$
    In particular, if $l\leq 2^{e}$, then $\sigma_{l,e^{\prime}}\sigma_{k,e}=\sigma_{l,e^{\prime}}$ if $l=k$, and 0 otherwise.
  4. (d)
    (Recursive split) $\sigma_{k,e}=\sigma_{k,e+1}+\sigma_{2^{e}+k,e+1}$.
  5. (e)
    Let $n=n_{0}+2n_{1}+\dots+2^{e-1}n_{e-1}$, $n_{i}=0,1$. Then, $\sigma_{k,e}=\frac{1}{2^{e}}\sum_{i=0}^{2^{e}-1}(-1)^{\sum i_{j}k_{j}}\lambda_{i}$.
  6. (f)
    $\sigma_{1,e}=\frac{1}{2^{e}}\delta_{2^{e}}=\frac{1}{2^{e}}\left(\sum_{j=0}^{2^{e}-1}(-1)^{j}\lambda_{j}\right)$.
  7. (g)
    $\alpha(\sigma_{k,e})=\frac{1}{2^{e}}$ for every $k$.
Proof.
Parts (f) and (g) follow directly from (e).
  1. (a)
    From Corollary 1.14 and the fact that $\lambda_{0}$ is the unit of the Han–Monsky ring, one checks that $(\lambda_{0}+\lambda_{2^{n}})^{2}=2(\lambda_{0}+\lambda_{2^{n}})$ and $(\lambda_{0}+\lambda_{2^{n}})(\lambda_{0}-\lambda_{2^{n}})=0.$ Applying the definition, the claim follows.
  2. (b)
    Follows directly from part (a).
  3. (c)
    Both by the fact that $k<2^{n}$ and $|l-k|\geq 2^{n}$ we conclude that $l_{j}=k_{j}$ for $j=0,\dots,n-1$. Thus,
    $\left(\prod^{n-1}_{j=0}(\lambda_{0}+(-1)^{k_{j}}\lambda_{2^{j}})\right)\cdot\left(\prod^{m-1}_{j=0}(\lambda_{0}+(-1)^{l_{j}}\lambda_{2^{j}})\right)=$
    $\left(\prod^{n-1}_{j=0}(\lambda_{0}+(-1)^{l_{j}}\lambda_{2^{j}})^{2}\right)\cdot\left(\prod^{m-1}_{j=n}(\lambda_{0}+(-1)^{l_{j}}\lambda_{2^{j}})\right)=2^{n}\prod^{m-1}_{j=0}(\lambda_{0}+(-1)^{l_{j}}\lambda_{2^{j}}).$
  4. (d)
    By definition, $\sigma_{k,e+1}+\sigma_{2^{e}+k,e+1}=\frac{1}{2}\left(\sigma_{k,e}\cdot(\lambda_{0}+\lambda_{2^{e}})+\sigma_{k,e}\cdot(\lambda_{0}-\lambda_{2^{e}})\right)=\sigma_{k,e}.$
  5. (e)
    By induction on $e$. The base case is clear since $\sigma_{0,1}=\frac{1}{2}(\lambda_{0}+\lambda_{1})$ and $\sigma_{1,1}=\frac{1}{2}(\lambda_{0}-\lambda_{1})$. Now, by the induction hypothesis,
    $\prod_{j=0}^{e-1}(\lambda_{0}+(-1)^{k_{j}}\lambda_{2^{j}})=(\lambda_{0}+(-1)^{k_{e-1}}\lambda_{2^{e-1}})\sum_{i=0}^{2^{e-1}-1}(-1)^{\sum_{j=0}^{e-2}i_{j}k_{j}}\lambda_{i},$
    which yields the formula.
    $\sigma_{1,e}:=\frac{1}{2^{e}}(\lambda_{0}-\lambda_{1})\prod_{j=1}^{e-1}(\lambda_{0}+\lambda_{2^{j}})=\frac{1}{2^{e}}\left(\lambda_{0}-\lambda_{1}+\dots+(-1)^{e-1}\lambda_{2^{e}-1}\right)=\frac{1}{2^{e}}\delta_{2^{e}}.$
Remark.
Since the $\sigma$-elements form a $\mathbb{Q}$-basis of $\Lambda_{e}\otimes\mathbb{Q}$, the statement of 3.2(a) represents the fact that $\Lambda_{e}\otimes\mathbb{Q}$ is isomorphic as a $\mathbb{Q}$-algebra to the direct product of $p^{e}$ copies of $\mathbb{Q}$, where the $\sigma$-elements would be the canonical basis of this product.
We provide two technical lemmas that we will need later:
Lemma 3.3.
Let $10\leq k\leq 2^{e}-$ 1, $e\geq 1$, and denote $k=k_{0}+2k_{1}+...+2^{e-1}k_{e-1}$ where $0\leq k_{j}\leq 1$ integers. Then, $\sigma_{k,e}\lambda_{2^{e}-1}=(-1)^{\sum_{j=0}^{e-1}k_{j}}\sigma_{k,e}.$
Proof.
By 3.2(e), recall that $\sigma_{k,e}=\sum_{i=0}^{2^{e}-1}(-1)^{\sum i_{j}k_{j}}\lambda_{i}$. Observe that $2^{e}-1=1+2+\dots+2^{e-1}$, so $2^{e}-i-1=(1-i_{0})+(1-i_{1})2+(1-i_{2})4+\dots+(1-i_{e-1})2^{e-1}$. Therefore,
$\lambda_{2^{e}-1}\sigma_{k,e}=\sum_{i=0}^{2^{e}-1}(-1)^{\sum i_{j}k_{j}}\lambda_{2^{e}-i-1}=\sum_{i=0}^{2^{e}-1}(-1)^{\sum(1-i_{j})k_{j}}\lambda_{i}=(-1)^{\sum_{j=0}^{e-1}k_{j}}\sigma_{k,e}.$
Lemma 3.4.
Let $10\leq k\leq 2^{e+1}-$ 1, $e\geq 1$. In the same notation as the previous lemma, we have that $\sigma_{k,e+1}\lambda_{2^{e}-1}=(-1)^{\sum_{j=0}^{e-1}k_{j}}\sigma_{k,e+1}.$
Proof.
Write $k^{\prime}=k$ if $k\leq 2^{e}-1$ and $k-2^{e}$ if $12^{e}\leq k\leq 2^{e+1}-$ 1. Then $\sigma_{k,e+1}=\sigma_{k^{\prime},e}(\lambda_{0}+(-1)^{k_{e}}\lambda_{2^{e}})$, and we apply Lemma 3.3:
$\lambda_{2^{e}-1}\sigma_{k,e+1}$ $=\lambda_{2^{e}-1}\sigma_{k^{\prime},e}(\lambda_{0}+(-1)^{k_{e}}\lambda_{2^{e}})$
$=(-1)^{\sum_{j=0}^{e-1}k_{j}}\sigma_{k^{\prime},e}(\lambda_{0}+(-1)^{k_{e}}\lambda_{2^{e}})=(-1)^{\sum_{j=0}^{e-1}k_{j}}\sigma_{k,e+1}.$
Remark.
Thus, for example, $\sigma_{2^{e},2^{e+1}}\lambda_{2^{e}-1}=\sigma_{2^{e}+1,2^{e+1}},$ and $\sigma_{2^{e},2^{e+1}}\lambda_{2^{e}-1}=-\sigma_{2^{e}+1,2^{e+1}}$.
3.2. $A_{1}$ singularity in characteristic 2
We will compute the Hilbert-Kunz function of $R_{2,d}$. When $d$ is even, Yoshida [Yos09] computed it and conjectured the formula when $d$ is odd. First, a lemma about an object of the Han–Monsky ring:
Lemma 3.5.
Let $k$ be a field of characteristic 2. For $e\geq 1$, set $\mu_{e}:=\left(k[x,y]/(x^{2^{e}},y^{2^{e}}),xy\right)$. Then $\mu_{e}=\delta_{2^{e}}+2\delta_{1}+\dots+2\delta_{2^{e}-1}.$
Proof.
Given that the $T$-action is given by product by a monomial, it is easy to see that the orbits of the monomials $1,x,...,x^{2^{n}-1},y,...,y^{2^{n}-1}$ under the action of $T$ produce disjoint $k$-vector subspaces that span the $k$-object – more specifically, the orbits are the following:
$\begin{array}[]{rllllll}1&\mapsto xy&\mapsto\dots&\mapsto x^{2^{e}-1}y^{2^{e}-1}&\mapsto 0,\\[5.0pt] x^{i}&\mapsto x^{i+1}y&\mapsto\dots&\mapsto x^{2^{e}-1}y^{2^{e}-i-1}&\mapsto 0,\\[5.0pt] y^{j}&\mapsto xy^{j+1}&\mapsto\dots&\mapsto x^{2^{e}-j-1}y^{2^{e}-1}&\mapsto 0.\end{array}$
According to this decomposition, one sees that $x^{i}$ (or $y^{i}$) generates a cyclic $k[T]$-submodule whose isomorphism class is $\delta_{2^{e}-i}$. ∎
Proposition 3.6.
Let $k$ be a field of characteristic 2. Then,
$\left(\frac{k[[x_{0},\dots,x_{d}]]}{(x_{0}^{[2^{e}]},\dots,x_{d}^{[2^{e}]})},Q_{d}\right)=\begin{cases}\mu_{e}^{m}\cdot 2\delta_{2^{e-1}}&(d=2m)\\ \mu_{e}^{m}&(d=2m-1)\end{cases}$
where $Q_{d}$ is as in Definition 2.1 and $\mu_{e}$ as in the previous lemma.
Proof.
If $d=2m$, $m\geq 1$, then $Q_{d}=x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{2}$. This implies that the $k$-object in question decomposes as:
$\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}\otimes\stackrel{{\scriptstyle(m)}}{{\dots}}\otimes\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}\otimes\frac{k[x]}{(x^{2^{e}})},$
where the $k$-object $\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}$ has $T$-action $xy$, the $k$-object $\frac{k[x]}{(x^{[2^{e}]})}$ has the $T$-action $y^{2}$.
If $d=2m-1$, $m\geq 1$, then $Q_{d}=x_{0}x_{1}+\dots+x_{d-3}x_{d-2}+x_{d-1}x_{d}$ and the $k$-object decomposes as
$\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}\otimes\stackrel{{\scriptstyle(m)}}{{\dots}}\otimes\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})},$
where the $k$-object $\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}$ has $T$-action $xy$. ∎
The aim of the following lemmas is expressing $\mu_{e}$ over the $\sigma$-basis:
Lemma 3.7.
For $e\geq 1$, $\mu_{e}=\sum_{k=0}^{2^{e}-1}(-1)^{k}(2(2^{e}-k)-1)\lambda_{k}$.
Proof.
We use the change of variables in the remark after Definition 1.7 on the expression in Lemma 3.5,
$2\delta_{1}+...+2\delta_{2^{e}-1}+\delta_{2^{e}}$ $=2\lambda_{0}+2(\lambda_{0}-\lambda_{1})+\cdots+2(\lambda_{0}-\dots+\lambda_{2^{e}-2})+(\lambda_{0}-\dots-\lambda_{2^{e}-1})$
$=\left[2\cdot(2^{e}-1)\lambda_{0}-2\cdot(2^{e}-2)\lambda_{1}+...+2\lambda_{2^{e}-2}\right]+\left[\lambda_{0}-\dots-\lambda_{2^{e}-1}\right]$
$=\sum_{k=0}^{2^{e}-1}(-1)^{k}(2(2^{e}-k-1)+1)\lambda_{k}.$
Corollary 3.8.
The expression $\mu_{e}$ verifies the following recursive formula:
$\mu_{e}=(\lambda_{0}-\lambda_{2^{e-1}})\mu_{e-1}+2^{2e-1}\sigma_{1,e-1}$
for every $e\geq 2$.
Proof.
We use the expression given by Lemma 3.7,
$\mu_{e}$ $=\sum_{k=0}^{2^{e}-1}(-1)^{k}(2(2^{e}-k)-1)\lambda_{k}$
$=\sum_{k=0}^{2^{e-1}-1}(-1)^{k}(2(2^{e}-k)-1)\lambda_{k}+\sum_{k=2^{e-1}}^{2^{e}-1}(-1)^{k}(2(2^{e}-k)-1)\lambda_{k}.$
The second summand is $-\lambda_{2^{e-1}}\mu_{e-1}$ by 1.12; as for the first summand,
$\sum_{k=0}^{2^{e-1}-1}(-1)^{k}(2(2^{e}-k)-1)\lambda_{k}$ $=\sum_{k=0}^{2^{e-1}-1}\left[(-1)^{k}(2(2^{e-1}-k)-1)+(-1)^{k}2\cdot 2^{e-1}\right]\lambda_{k}=$
$=\sum_{k=0}^{2^{e-1}-1}(-1)^{k}(2(2^{e-1}-k)-1)\lambda_{k}+2^{e}\sum_{k=0}^{2^{e-1}-1}(-1)^{k}\lambda_{k}=$
$=\mu_{e-1}+2^{e}2^{e-1}\sigma_{1,e-1}.$
and the result follows from the same lemma and 3.2(f).
Alternatively, one can deduce the same recursive formula from Lemma 3.5 directly observing that $(\mu_{e}-\delta_{2^{e}})=(\lambda_{0}+\lambda_{2^{e-1}})(\mu_{e-1}-\delta_{2^{e-1}})+2^{e-1}\delta_{2^{e-1}}$, using the definition of $\lambda_{i}$ and Lemma 3.3 in [HM93]. ∎
Remark.
Using the notation and results in [MT06], one can deduce the results above in a way that is similar in essence. Using the formula for $\varphi_{fg}$ provided in Proposition 4.2 of that paper, one directly obtains that $\varphi:=\varphi_{xy}=2t-t^{2}$, which is immediately seen to satisfy the functional equations $4\varphi(t/2)=\varphi+t$ and $34\varphi((t+1)/2)=\varphi+$ 3. Using arguments very similar to those in Example 1, one concludes that $4S(a)=a+\lambda_{1}a+2(\lambda_{0}-\lambda_{1})\Delta$, where $a=\mathscr{L}(\varphi)$. It is not difficult to see that $(\lambda_{0}-\lambda_{1})\Delta=2S(\Delta)$, which shows that the formula obtained is equivalent to that in Corollary 3.8.
Once we have these elements, using the recursive formula for $\mu_{e}$, the computation of the Hilbert-Kunz function is not too difficult.
Proposition 3.9.
Given $e\geq 1$, then $\mu_{e}=2^{e}(\sigma_{0,e}+2^{e}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{k}\sigma_{2^{k}+1,e}).$
Proof.
We prove it by induction on $e$. By Proposition 3.8, $M_{1}=3\lambda_{0}-\lambda_{1}=2\sigma_{0,1}+4\sigma_{1,1},$ so the result holds when $e=1$ (where we regard $\sum_{k=1}^{0}2^{k}\sigma_{2^{k}+1,1}=0$.)
To show the general $e\geq 1$ case, we will employ the recursive formula in Proposition 3.8, the property that $(\lambda_{0}+\lambda_{2^{e}})\sigma_{k,e}=2\sigma_{k,e+1}$, and Proposition 3.2(d):
$\mu_{e+1}$ $=(\lambda_{0}+\lambda_{2^{e}})\mu_{e}+2^{2(e+1)-1}\sigma_{1,e}$
$=2^{e}(\lambda_{0}+\lambda_{2^{e}})(\sigma_{0,e}+2^{e}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{k}\sigma_{2^{k}+1,e})+2^{2e+1}\sigma_{1,e}$
$=2^{e}\cdot 2(\sigma_{0,e+1}+2^{e}\sigma_{1,e+1}+\sum_{k=1}^{e-1}2^{k}\sigma_{2^{k}+1,e+1})+2^{2e+1}(\sigma_{1,e+1}+\sigma_{2^{e}+1,e+1})$
$=2^{e+1}(\sigma_{0,e+1}+2^{e}\sigma_{1,e+1}+\sum_{k=1}^{e-1}2^{k}\sigma_{2^{k}+1,e+1}+2^{e}\sigma_{1,e+1}+2^{e}\sigma_{2^{e}+1,e+1}).$
The result follows from the last line. ∎
Theorem 3.10.
For $d\geq 2$,
$\text{HK}_{e}(R_{2,d})=\begin{dcases}\frac{2^{m}}{2^{m}-1}2^{de}-\frac{(2^{m-1})^{e}}{2^{m}-1},&\text{ if }d=2m-1;\\ \frac{2^{m}+1}{2^{m}}2^{de},&\text{ if }d=2m.\end{dcases}$
where $R_{p,d}$ is as in Definition 2.1. In particular, the Hilbert–Kunz multiplicity of an $A_{1}$ singularity in characteristic 2 is
$e_{\text{HK}}\left(R_{2,d}\right)=\begin{dcases}\frac{2^{m}}{2^{m}-1},&\text{ if }d=2m-1;\\ \frac{2^{m}+1}{2^{m}},&\text{ if }d=2m.\\ \end{dcases}$
Proof.
Firstly, let $d=2m-1$. The result follows from Proposition 3.9, using idempotency and orthogonality of the $\sigma$-elements and Proposition 3.2(g):
$\text{HK}_{e}(R_{2,d})$ $=\alpha(\mu_{e}^{m})=2^{em}\alpha(\sigma_{0,e}+2^{em}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{km}\sigma_{2^{k}+1,e})$
$=2^{em}\sum_{k=0}^{e}2^{km}\frac{1}{2^{e}}=2^{e(m-1)}\sum_{k=0}^{e}2^{km}=\frac{2^{(e+1)m}-1}{2^{m}-1}.$
So the formula follows in the odd-dimensional case.
Now, for the even-dimensional case, $d=2m$, observe that $R_{2,d}/\mathfrak{m}^{[2^{e}]}=2\delta_{2^{e-1}}\mu_{e}^{m}$, and by Proposition 3.2(f), we have that $\delta_{2^{e-1}}=2^{e-1}\sigma_{1,e-1}=2^{e-1}(\sigma_{1,e}+\sigma_{2^{e-1}+1,e}).$ Then, we proceed similarly – by Proposition 3.6, we have that $\text{HK}_{e}(R_{2,2m})=\alpha(\mu_{e}^{m}\cdot 2\delta_{2^{e-1}}),$ and
$\mu_{e}^{m}\cdot 2\delta_{2^{e-1}}$ $=2^{em}(\sigma_{0,e}+2^{em}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{km}\sigma_{2^{k}+1,e})\cdot 2^{e}(\sigma_{1,e}+\sigma_{2^{e-1}+1,e})$
$=2^{e(m+1)}(2^{em}\sigma_{1,e}+2^{(e-1)m}\sigma_{2^{e-1}+1,e}).$
Again, using that $\alpha(\sigma_{k,e})=\frac{1}{2^{e}}$, from applying $\alpha$ to the above, we obtain that $\text{HK}_{n}(R_{2,d})$ is
$2^{e(m+1)}(2^{em}\frac{1}{2^{e}}+2^{(e-1)m}\frac{1}{2^{e}})=2^{em}(2^{em}+2^{(e-1)m})=2^{nd}+\frac{1}{2^{m}}2^{dn}=\frac{2^{m}+1}{2^{m}}2^{dn}.$
3.3. $A_{2}$ singularity in characteristic 2
Proposition 3.11.
Let $k$ be a field of characteristic 2. Then, in the Han–Monsky ring
$\left(\frac{k[[x_{0},\dots,x_{d}]]}{(x_{0}^{[2^{e}]},\dots,x_{d}^{[2^{e}]})},P_{d}\right)=\begin{cases}\mu_{e}^{m}3\delta_{2^{e}/3}&(d=2m)\\ \mu_{e}^{m-1}2\delta_{2^{e-1}}3\delta_{2^{e}/3}&(d=2m-1)\end{cases}$
where $P_{d}$ was defined in Definition 2.1, and $\mu_{e}$ is as defined in Lemma 3.5.
Proof.
Let $d=2m+1$. Then, $P_{d}=x_{0}x_{1}+\dots+x_{d-3}x_{d-2}+x_{d-1}^{2}+x_{d}^{3}$ and the object decomposes as
$\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}\otimes\dots\otimes\frac{k[x,y]}{(x^{[2^{e}]},y^{[2^{e}]})}\otimes\frac{k[y]}{(y^{[2^{e}]})}\otimes\frac{k[x]}{(x^{[2^{e}]})},$
where the $k$-object $k[x,y]/(x^{[2^{e}]},y^{[2^{e}]})$ has $T$-action $xy$, the $k$-object $k[y]/(y^{[2^{e}]})$ has the $T$-action $y^{2}$, and finally the $k$-object $k[x]/(x^{[2^{e}]})$ has $T$-action $x^{3}$. We already know that the first is $\mu_{e}$ and the second $2\delta_{2^{e-1}}$. The third $k$-object is $3\delta_{2^{e}/3}$, as one deduces from Lemma 2.2 for $s=0$.
If $d=2m$, then $P_{d}=x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{3}$ and the result follows from the same discussion. ∎
Lemma 3.12.
For $e\geq 1$, let $\xi_{e}:=3\delta_{2^{e}/3}$. Then, $\xi_{1}=2\lambda_{0}=2\sigma_{0,1}+2\sigma_{1,1},\xi_{2}=3\lambda_{0}-\lambda_{1}=2\sigma_{0,2}+4\sigma_{1,2}+2\sigma_{2,2}+4\sigma_{3,2}$,
$\xi_{e+1}-\lambda_{2^{e}-1}\xi_{e}=3\delta_{2^{e}},e\geq 1.$
Proof.
Let $e\geq 1$, and denote $2^{e}=3d+r$ and $2^{e+1}=3D+R$, where $d,r,D,R\in\mathbb{Z}$ and $30\leq r,R<$ 3. Then, from Definition 1.15,
$\xi_{e+1}$ $=3\lambda_{0}-3\lambda_{1}+\dots+(-1)^{D-1}3\lambda_{D-1}+(-1)^{D}R\lambda_{D},$
$\xi_{e}$ $=3\lambda_{0}-3\lambda_{1}+\dots+(-1)^{D-1}3\lambda_{d-1}+(-1)^{d}r\lambda_{d}.$
Then, using 1.12(b) one obtains
$\xi_{e+1}-\lambda_{2^{e}-1}\xi_{e}=(3\lambda_{0}-\dots+(-1)^{D}R\lambda_{D})-\\ -((-1)^{d}r\lambda_{2^{e}-d-1}+(-1)^{d-1}3\lambda_{2^{e}-d-2}+\dots-3\lambda_{2^{e}-2}+3\lambda_{2^{e}-1}).$ (3)
One can check that $D=2^{e}-(d+1)$ and $R+r=3$, which implies that $(-1)^{D}R\lambda_{D}-(-1)^{d}r\lambda_{2^{e}-d-1}=(-1)^{D}3\lambda_{D}$, which implies that the RHS in (3) is $3\delta_{2^{e}}$. ∎
The point of the following lemma is giving a formula of the coefficients of $\xi_{e}$ over the $\sigma$-basis that we will need since the rest of them get canceled in the computation of the Hilbert–Kunz function of the $A_{2}$ singularity, as we will see in Theorem 3.14.
Lemma 3.13.
Let $\xi_{e}$ as in Lemma 3.12. Let $a_{k,e}\in\mathbb{Q}$ such that $\xi_{e}=\sum_{k=0}^{2^{e}-1}a_{k,e}\sigma_{k,e}$. Then $a_{0,e}=2$, $a_{1,e}=2^{e+1}$, $a_{2^{k}+1,e}=2^{k+1}$ for $k=1,\dots,e-2$, $e\geq 2$.
Proof.
Throughout the proof, we use that $\sigma_{k,e}\xi_{e}=a_{k,e}\sigma_{k,e}$, and the rest of the orthogonality and idempotency properties (3.2(a), (b) and (c)) without further mention. Also, $e\geq 2$.
  1. (1)
    We start computing $a_{0,e}$ and $a_{1,e}$, for $e\geq 2$. Since $\delta_{2^{e}}=2^{e}\sigma_{1,e}$ by 3.2(f), it follows that $\sigma_{0,e+1}\delta_{2^{e}}=0$, and $\sigma_{1,e+1}\delta_{2^{e}}=2^{e}\sigma_{1,e+1}$. Also, by Lemma 3.4, $\sigma_{0,e+1}\lambda_{2^{e}-1}=\sigma_{0,e+1}$ and $\sigma_{1,e+1}\lambda_{2^{e}-1}=-\sigma_{1,e+1}$. Thus, multiplying the recurrence in Lemma 3.12 by $\sigma_{0,e+1}$ and $\sigma_{1,e+1}$, we obtain in each case
    $\sigma_{0,e+1}\xi_{e+1}-\sigma_{0,e+1}\lambda_{2^{e}-1}\xi_{e}=0$ $\Longrightarrow a_{0,e+1}-a_{0,e}=0,$
    $\sigma_{1,e+1}\xi_{e+1}-\sigma_{1,e+1}\lambda_{2^{e}-1}\xi_{e}=3\cdot 2^{e}\sigma_{1,e+1}$ $\Longrightarrow a_{1,e+1}+a_{1,e}=3\cdot 2^{e}.$
    Given that $a_{0,1}=a_{1,1}=2$ by 3.12, we have that $a_{0,e}=2$ and $a_{1,e}=2^{e}$ for $e\geq 1$.
  2. (2)
    In order to compute $a_{2^{k}+1,e}$ for $21\leq k\leq e-$ 2, we first need to compute $a_{2^{e-1}+1,e}$ for every $e\geq 2$. Using the recurrence in 3.12 again,
    $a_{2^{e-1}+1,e}\sigma_{2^{e-1}+1,e}=\sigma_{2^{e-1}+1,e}\xi_{e}=\lambda_{2^{e-1}-1}\sigma_{2^{e-1}+1,e}\xi_{e-1}+3\sigma_{2^{e-1}+1,e}\delta_{2^{e-1}}.$ (4)
    As for the product $\sigma_{2^{e-1}+1,e}\delta_{2^{e-1}}$ above, by 3.2(f), we have that $\sigma_{2^{e-1}+1,e}\delta_{2^{e-1}}=\sigma_{2^{e-1}+1,e}2^{e-1}\sigma_{1,e-1}=2^{e-1}\sigma_{2^{e-1}+1,e},$ using that $\sigma_{2^{e-1}+1,e}\sigma_{1,e-1}=\sigma_{2^{e-1}+1,e}$.
    About the first summand in (4), observe that for $e\geq 2$, $\lambda_{2^{e-1}-1}\sigma_{2^{e-1}+1,e}=\sigma_{2^{e-1}+1,e}$ by Lemma 3.4, so we continue with the RHS of (4):
    $3\sigma_{2^{e-1}+1,e}\delta_{2^{e-1}}-\sigma_{2^{e-1}+1,e}\lambda_{2^{e-1}-1}\xi_{e-1}=3\cdot 2^{e-1}\sigma_{2^{e-1}+1,e}-\sigma_{2^{e-1}+1,e}\xi_{e-1}.$
    Finally, from 3.2(d),
    $\sigma_{2^{e-1}+1,e}\xi_{e-1}$ $=\sigma_{2^{e-1}+1,e}\sum a_{s,e-1}\sigma_{s,e-1}$
    $=\sigma_{2^{e-1}+1,e}\left(\sum_{s=0}^{2^{e-1}-1}a_{s,e-1}\sigma_{s,e}+\sum_{s=2^{e-1}}^{2^{e}-1}a_{s,e-1}\sigma_{2^{e-1}+s,e}\right)=a_{1,e-1}\sigma_{2^{e-1}+1,e}.$
    We conclude from this and (4) that $a_{2^{e-1}+1,e}=3\cdot 2^{e-1}-a_{1,e-1}=2\cdot 2^{e-1}=2^{e}$, and therefore,
    $a_{2^{e-1}+1,e}=2^{e},e\geq 2.$ (5)
  3. (3)
    We now compute $a_{2^{k}+1,e}$ for $21\leq k\leq e-$ 2 and $e\geq 2$. Note that, for those $k$ and $e$, it happens again that $\sigma_{2^{k}+1,e+1}\delta_{2^{e}}=0$, so multiplying the recurrence in 3.12 by $\sigma_{2^{k}+1,e+1}$, we obtain
    $\sigma_{2^{k}+1,e+1}\xi_{e+1}-\sigma_{2^{k}+1,e+1}\lambda_{2^{e}-1}\xi_{e}=0\Longrightarrow a_{2^{k}+1,e+1}-a_{2^{k}+1,e}=0.$
    This implies that for all $21\leq k\leq e-$ 2, $a_{2^{k}+1,e}=a_{2^{k}+1,k+1}$, and we know by the previous part (setting $e=k+1$ in (5)) that $a_{2^{k}+1,k+1}=2^{k+1}$ for $k\geq 2$. Therefore, we conclude that
    $a_{2^{k}+1,e}=2^{k+1},1\leq k\leq e-2,e\geq 2.$
Table 2 summarizes the results in the previous lemmas.
$a_{0,e}$ $a_{1,e}$ $a_{2^{k}+1,e},1\leq k\leq e-1$
$2$ $2^{e}$ $2^{k+1}$
Table 2. Critical coefficients of $x^{3}$.
Theorem 3.14.
The Hilbert–Kunz function of an $A_{2}$ singularity in characteristic 2 is
$\text{HK}_{e}(S_{2,d})=\begin{dcases}\frac{2^{m+1}+1}{2^{m+1}-1}2^{de}-\frac{2}{2^{m+1}-1}2^{e(m-1)}&(d=2m);\\ \frac{2^{m}+1}{2^{m}}2^{de}&(d=2m-1);\end{dcases}$
with $S_{p,d}$ as in Definition 2.1. In particular, the Hilbert–Kunz multiplicity of an $A_{2}$ singularity in characteristic 2 is
$e_{\text{HK}}\left(S_{2,d}\right)=\begin{dcases}\frac{2^{m+1}+1}{2^{m+1}-1}&(d=2m);\\ \frac{2^{m-1}+1}{2^{m-1}}&(d=2m-1);\\ \end{dcases}$
Proof.
We deal with even and odd-dimensional cases separately.
Firstly, let $d=2m+1$. By 3.11, $\text{HK}_{e}(S_{3,2m+1})=\alpha\left(\mu_{e}^{m}\xi_{e}\right)$. Let $\xi_{e}$ as in 3.13, and $\mu_{e}$ as in 3.9, so
$\mu_{e}^{m}\xi_{e}$ $=2^{em}(\sigma_{0,e}+2^{em}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{km}\sigma_{2^{k}+1,e})\sum_{k=0}^{2^{e}-1}a_{k,e}\sigma_{k,e}$
$=2^{em}(a_{0,e}\sigma_{0,e}+2^{em}a_{1,e}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{km}a_{2^{k}+1,e}\sigma_{2^{k}+1,e}).$
Using Lemma 3.13, the orthogonality properties of $\sigma$-basis, 3.2, and the linearity of $\alpha$,
$\alpha(\mu_{e}^{m}\xi_{e})$ $=2^{em}\alpha(a_{0,e}\sigma_{0,e}+2^{em}a_{1,e}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{km}a_{2^{k}+1,e}\sigma_{2^{k}+1,e})$
$=2^{em}(2\frac{1}{2^{e}}+2^{em}2^{e}\frac{1}{2^{e}}+\sum_{k=1}^{e-1}2^{km}2^{k+1}\frac{1}{2^{e}})$
$=2^{em}(2+2^{e(m+1)}+2\sum_{k=1}^{e-1}2^{k(m+1)})$
$=2^{em}(-2^{e(m+1)}+2\sum_{k=0}^{e}2^{k(m+1)})=2^{em}(-2^{e(m+1)}+2\frac{2^{(e+1)(m+1)}-1}{2^{m+1}-1}).$
Thus,
$\text{HK}_{e}(S_{2,d})=2^{em}(-2^{e(m+1)}+2\frac{2^{(e+1)(m+1)}-1}{2^{m+1}-1})=-2^{de}+2\frac{2^{m+1}2^{de}-2^{e(m-1)}}{2^{m+1}-1},$
from which the result follows for $d=2m+1$.
Let now $d=2m$. Recall that by Proposition 3.2, we have that $\delta_{2^{e-1}}=2^{e-1}\sigma_{1,e-1}=2^{e-1}(\sigma_{1,e}+\sigma_{2^{e-1}+1,e})$, so $2\delta_{2^{e-1}}=2^{e}(\sigma_{1,e}+\sigma_{2^{e-1}+1,e})$. Now, by Proposition 3.11, we have that $\text{HK}_{e}(S_{3,2m})=\alpha\left(2\mu_{e}^{m-1}\delta_{2^{e-1}}\xi_{e}\right)$, so
$\mu_{e}^{m-1}2\delta_{2^{e-1}}\xi_{e}$ $=2^{e(m-1)}(\sigma_{0,e}+2^{e(m-1)}\sigma_{1,e}+\sum_{k=1}^{e-1}2^{k(m-1)}\sigma_{2^{k}+1,e})2^{e}(\sigma_{1,e}+\sigma_{2^{e-1}+1,e})\sum_{k=0}^{2^{e}-1}a_{k,e}\sigma_{k,e}$
$=2^{em}2^{e(m-1)}2^{e}a_{1,e}\sigma_{1,e}+2^{em}2^{(e-1)(m-1)}2^{e}a_{2^{e-1}+1,e}\sigma_{2^{e-1}+1,e}.$
Thus, applying $\alpha$ to the above, and using Lemma 3.13,
$\text{HK}_{e}(S_{2,d})=2^{de}2^{e}\frac{1}{2^{e}}+2^{de}\frac{1}{2^{m}}2^{e}\frac{1}{2^{e}}=\frac{2^{m}+1}{2^{m}}2^{de}.$
3.4. $A_{1}$ and $A_{2}$ singularity in characteristic 3
Here, we compute the Hilbert–Kunz multiplicity of the $A_{2}$ singularity.
Theorem 3.15.
For $d\geq 2$,
$e_{\text{HK}}\left(R_{3,d}\right)=1+\frac{3\cdot 2^{d}}{3^{d+1}-2^{d}+(-1)^{d}}\quad\text{and}\quad e_{\text{HK}}\left(S_{3,d}\right)=1+\left(\frac{2}{3}\right)^{d-1}$
Proof.
The Hilbert–Kunz multiplicity of $R_{3,d}$ was already computed in [Yos09].
As for $S_{3,d}$, first note that $k[[x_{0},\dots,x_{d-1},x_{d}^{3}]]\subset k[[x_{0},\dots,x_{d}]]$ is a free extension of rank $3$. Thus, the following chain of inequalities follows:
$\text{HK}_{e}(S_{3,d})$ $=\dim_{k}\frac{k[[x_{0},\dots,x_{d}]]}{(x_{0}^{3^{e}},\dots,x_{d}^{3^{e}})+(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3})}$
$=3\cdot\dim_{k}\frac{k[[x_{0},\dots,x_{d}^{3}]]}{(x_{0}^{3^{e}},\dots,x_{d}^{3^{e}})+(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3})}$
$=3\cdot\dim_{k}\frac{k[[x_{0},\dots,x_{d-1}]]}{(x_{0}^{3^{e}},\dots,x_{d-1}^{3^{e}},(x_{0}^{2}+\dots+x_{d-1}^{2})^{3^{e-1}})},$
and now, since $k[[x_{0}^{3^{e-1}},\dots,x_{d-1}^{3^{e-1}}]]\subset k[[x_{0},\dots,x_{d-1}]]$ is also a free extension of rank $3^{d(e-1)}$,
$\text{HK}_{e}(S_{3,d})$ $=3\cdot 3^{(e-1)d}\cdot\dim_{k}\frac{k[[x_{0}^{3^{e-1}},\dots,x_{d-1}^{3^{e-1}}]]}{(x_{0}^{3^{e}},\dots,(x_{0}^{2}+\dots+x_{d-1}^{2})^{3^{e-1}})}$
$=3\cdot 3^{(e-1)d}\cdot\dim_{k}\frac{k[[y_{0},\dots,y_{d-1}]]}{(y_{0}^{3},\dots,y_{d-1}^{3})+(y_{0}^{2}+\dots+y_{d-1}^{2})}=3\cdot 3^{(e-1)d}\cdot\text{HK}_{1}(R_{3,d-1}).$
In the mentioned Yoshida’s notes, it is proven that $\text{HK}_{1}(R_{3,d-1})=2^{d-1}+3^{d-1}$, so
$e_{\text{HK}}\left(S_{3,d}\right)=\frac{3\cdot 3^{d(e-1)}\cdot(2^{d-1}+3^{d-1})}{3^{de}}=1+\left(\frac{2}{3}\right)^{d-1}$
A corollary to the formula of $e_{\text{HK}}\left(S_{3,d}\right)$ is that the $A_{2}$ singularity in characteristic 3 is a counterexample to Conjecture 2.7 in [JNS+23]:
Corollary 3.16.
$e_{\text{HK}}\left(S_{3,d}\right)<e_{\text{HK}}\left(R_{2,d}\right)$ if and only if $d\geq 8$ even and $d\geq 13$ odd.
Proof.
It follows from direct computation with the formula in Theorems 3.15 and 3.10. ∎
3.5. The inequality.
By the results in the previous chapter of this paper, it easily follows that the Hilbert–Kunz multiplicity of an $A_{2}$ singularity and an $A_{1}$ singularity are different:
Theorem 3.17 (Main theorem in characteristic 2).
For $d\geq 2$, $e_{\text{HK}}\left(S_{2,d}\right)>e_{\text{HK}}\left(R_{2,d}\right)$.
Proof.
Let $d=2m-1$. Then, from Theorems 3.10 and 3.14,
$e_{\text{HK}}\left(R_{2,d}\right)<e_{\text{HK}}\left(S_{2,d}\right)\Leftrightarrow\frac{2^{m}}{2^{m}-1}<\frac{2^{m-1}+1}{2^{m-1}} \Leftrightarrow 2^{d}<2^{d}+2^{m-1}-1,$
which is true for every $m>1$. If $d=2m$, from loc. cit.
$e_{\text{HK}}\left(R_{2,d}\right)<e_{\text{HK}}\left(S_{2,d}\right)\Leftrightarrow\frac{2^{m}+1}{2^{m}}<\frac{2^{m+1}+1}{2^{m+1}-1} \Leftrightarrow 2\cdot 2^{d}+2^{m}-1<2\cdot 2^{d}+2^{m},$
which is again true for every $m\geq 1$. ∎
Theorem 3.18 (Main theorem in characteristic 3).
For $d\geq 2$, $e_{\text{HK}}\left(S_{3,d}\right)>e_{\text{HK}}\left(R_{3,d}\right)$.
Proof.
By Theorem 3.15, we have to compare the terms $(2/3)^{d-1}$ and $3\cdot 2^{d}/(3^{d+1}-2^{d}+(-1)^{d})$. Now,
$(2/3)^{d-1}>3\cdot 2^{d}/(3^{d+1}-2^{d}+(-1)^{d})\Leftrightarrow 3^{d+1}-2^{d}+(-1)^{d}>3^{d}\cdot 2\Leftrightarrow 3^{d}-2^{d}+(-1)^{d}>0,$
where the second and last equivalence hold because $03^{d}-2^{d}+(-1)^{d}>$ 0 for all $d\geq 2$. ∎
Corollary 3.19.
Let $(R,\mathfrak{m})$ be a non-regular complete intersection of characteristic $p>3$ and dimension $d\geq 2$. Then, $e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(R_{p,d}\right).$ Moreover, if $\hat{R}\not\cong R_{p,d}$, then
$e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right).$ (6)
In particular, the strong Watanabe–Yoshida conjecture holds for complete intersections of characteristic $p>3$.
4. Strong Watanabe–Yoshida conjecture for complete intersections of characteristic 2 and 3.
[…]
Notation.
We will use multi-index notation in the following lemmas, so we will write $f=\sum f_{\underline{e}}x_{1}^{e_{1}}\dots x_{n}^{e_{n}}$ where ${\underline{e}}:=(e_{1},\dots,e_{n})$, and we define $|{\underline{e}}|=\sum e_{i}$ the norm of the multi-index.
4.1. Characteristic 2 case.
Lemma 4.1.
Let $k=\bar{k}$ be a field of characteristic 2 and $v,w\in k[[x_{0},\dots,x_{d}]]$, where $v$ is a unit. Then,
  1. (a)
    There exists $t\in k[[x_{0},\dots,x_{d}]]$ such that $t^{2}+vt+w=0$. Moreover, if $w$ is a unit, then $t$ also is.
  2. (b)
    There exists a unit $s\in k[[x_{0},\dots,x_{d}]]$ such that $s^{3}=v$.
Proof.
We first prove (a). Let $u=v^{-1}$. A solution to the equation exists if and only if $t$ exists such that $t=v^{-1}(t^{2}+w)=u(w+t^{2})$, which is the case if and only if the following set of equations has a solution over $k$:
$t_{{\underline{e}}}=\sum_{{\underline{e}}^{\prime}+{\underline{e}}^{\prime\prime}={\underline{e}}}u_{{\underline{e}}^{\prime}}(w_{{\underline{e}}^{\prime\prime}}+\sum_{{\underline{f}}+{\underline{f}}^{\prime}={\underline{e}}^{\prime\prime}}t_{{\underline{f}}}t_{{\underline{f}}^{\prime}}).$ (7)
Observe that $t_{\underline{e}}$ only appears on the RHS only if ${\underline{f}}=0,{\underline{f}}^{\prime}={\underline{e}}$ or ${\underline{f}}={\underline{e}}$ and ${\underline{f}}^{\prime}=0$. In both these cases, we have that $t_{{\underline{f}}}t_{{\underline{f}}^{\prime}}=t_{0}t_{{\underline{e}}}.$ Because $k$ has characteristic 2, these two terms cancel each other out. Thus, for every ${\underline{e}}$, the expression on the right side depends on coefficients of $t$ of monomials of strictly lower degree, i.e., it depends on coefficients $t_{{\underline{f}}},t_{{\underline{f}}^{\prime}}$ such that $|{\underline{f}}|,|{\underline{f}}^{\prime}|<|{\underline{e}}|.$
The equation for ${\underline{e}}=0$ has a solution because the field is algebraically closed, so it follows by induction on the norm of the multi-index that the set of equations (7) has a solution.
Let us now prove (b). Let $u=\sum b_{\underline{e}}x_{0}^{e_{0}}\dots x_{d}^{e_{d}}$. There exists a series $s=\sum a_{\underline{e}}x_{0}^{e_{0}}\dots x_{d}^{e_{d}}$ such that $s^{3}=u$ if and only if the following systems of equations have a solution:
$\sum_{\underline{i}+\underline{j}+\underline{k}={\underline{e}}}a_{\underline{i}}a_{\underline{j}}a_{\underline{k}}=b_{{\underline{e}}},{\underline{e}}\in\mathbb{Z}^{d+1}_{\geq 0}.$ (8)
The fact that these systems of equations have a solution in $k$ follows from induction on the norm of the multi-index ${\underline{e}}$. ∎
The following lemma is the characteristic 2 analog of Lemma 3.1 in [ES05], used for completing squares in characteristic not 2.
Lemma 4.2 (Completing squares in characteristic 2).
Let $k$ be an algebraically closed field of characteristic 2 and put $A=k[[x_{2},\dots,x_{d}]]$. Consider $B=A[[x_{0},x_{1}]]$ and $F=x_{0}x_{1}+Q(x_{2},\dots,x_{d})+G$ with $Q\in A$ a quadratic form and $G\in\mathfrak{m}_{B}^{3}$, where $\mathfrak{m}_{B}$ is the maximal ideal of $B$. Then there exist $a_{0},a_{1},b_{0},b_{1}\in B$ units, $a_{2},b_{2}\in(x_{2},\dots,x_{d})B$ and $G_{1}\in(x_{2},\dots,x_{d})^{3}B$ such that
$F=(a_{0}x_{0}+a_{1}x_{1}+a_{2})(b_{0}x_{0}+b_{1}x_{1}+b_{2})+Q(x_{2},\dots,x_{d})+G_{1}.$
Proof.
Let $x=x_{0}$ and $y=x_{1}$ for simplicity. We can assume that $F=x^{2}+xy+y^{2}+Q_{d-2}(x_{2},\dots,x_{d})$ after a change of variables. Then, note that we can express $G$ in the following way:
$G=c+ax+by+Ax^{2}+Bxy+Cy^{2},$ (9)
where $c\in(x_{2},\dots,x_{d})^{3},$ $a,b\in(x_{2},\dots,x_{d})^{2},$ $A,B,C\in(x_{2},\dots,x_{d}).$
Let $u:=1+A$, $v:=1+B$ and $w:=1+C$, which are units. Due to Lemma 4.1, there exists a unit $t$ such that $ut^{2}+vt+w=0$. The result follows by letting
$\begin{array}[]{lll}a_{0}=u,&a_{1}=t^{-1}w,&a_{2}=uv^{-1}b+t^{-1}wv^{-1}a,\\ b_{0}=1,&b_{1}=t,&b_{2}=v^{-1}ta+v^{-1}b,\end{array}$
and $G_{1}=c+a_{2}b_{2}$. ∎
Lemma 4.3.
Using the notation above, there is a change of variables transforming $(a_{0}x+a_{1}y+a_{2})(b_{0}x+b_{1}y+b_{2})$ into $xy.$
Proof.
Let us define a change of variables where $x_{i}\mapsto x_{i}$ for $i\geq 2$, and
$\left(\begin{array}[]{c}x\\ y\end{array}\right)\mapsto\left(\begin{array}[]{cc}a_{0}&a_{1}\\ b_{0}&b_{1}\end{array}\right)^{-1}\left(\begin{array}[]{c}x\\ y\end{array}\right).$
This matrix is invertible because its determinant is $a_{0}b_{1}+a_{1}b_{0}=v,$ which is a unit. ∎
We will also use the following lemma of quadratic forms in characteristic 2 in the main result of the section.
Corollary 4.4 (Diagonalization of quadratic forms in characteristic 2).
Let $k$ be an algebraically closed field of characteristic 2. Let $Q$ be a quadratic form over $k$. Then, $Q=Q_{k}$ for some $k$, for $Q_{k}$ as in Definition 2.1.
Proof.
See Theorem 7.31 in [EKM08]. ∎
The following result is the extension of Theorem 3.4 in [ES05] to characteristic 2.
Theorem 4.5.
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional singular complete local hypersurface ring of characteristic $p=2$, where $k$ is algebraically closed. Then one of the following is true:
  1. (1)
    $R\cong k[[x_{0},\dots,x_{d}]]/(Q_{d}),$ or
  2. (2)
    $e_{HK}(R)\geq e_{HK}(k[[x_{0},\dots,x_{d}]]/(P_{d}))$,
where $P_{d}$ is as defined in Definition 2.1.
Proof.
Set $R=k[[x_{0},...,x_{d}]]/(F),$ where we write $F=Q+G$ for $Q$ the homogeneous part of degree $2$ and $G\in\mathfrak{m}^{3}$. If $R\ncong R_{d,2},$ we will prove that $e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)$. Thanks to Lemma 4.4, we can assume, after a suitable change of variables, that $Q=Q_{k}$ for some $k<d$, where $Q_{k}$ is as in Definition 2.1 (note that the inequality is strict). Let us now define the polynomial with which we will deform the hypersurface. Let $\delta=0$ if the coefficient of $x_{d}^{3}$ in $G$ is non-zero, and $\delta=1$, otherwise.
  1. (1)
    If $k$ odd, then define
    $g:=\begin{cases}x_{k+1}x_{k+2}+\dots+x_{d-1}^{2}+\delta x_{d}^{3}&\text{if $d$is odd},\\ x_{k+1}x_{k+2}+\dots+x_{d-2}x_{d-1}+\delta x_{d}^{3}&\text{if $d$ is even}.\end{cases}$
  2. (2)
    If $k$ is even, then let
    $g=\begin{cases}x_{k}x_{k+1}+\dots+x_{d-1}^{2}+\delta x_{d}^{3}&\text{if $d$ isodd},\\ x_{k}x_{k+1}+\dots+x_{d-2}x_{d-1}+\delta x_{d}^{3}&\text{if $d$ is even}.\end{cases}$
Due to the upper semi-continuity of the Hilbert-Kunz multiplicity, Theorem 2.6 in [ES05], there exists a dense $\Lambda\subset k$ such that $\forall\alpha\in\Lambda\setminus\{0\},$ then
$e_{HK}\left(k[[x_{0},\dots,x_{d}]]/(F+\alpha g)\right)\leq e_{HK}\left(k[[x_{0},\dots,x_{d}]]/(F)\right).$
We make a different transformation depending on the parity of $k$:
  1. (1)
    If $k$ is odd, then $F_{\alpha}=F+\alpha g$ takes the following expression
    $x_{0}x_{1}+\dots+x_{k-1}x_{k}+\alpha x_{k+1}x_{k+2}+\dots+\begin{cases}\alpha x_{d-1}^{2}+\alpha\delta x_{d}^{3}+G&\text{if $d$ is odd},\\ \alpha\delta x_{d}^{3}+G&\text{if $d$ is even}.\end{cases}$
    Re-scaling the coordinates we can assume that $\alpha=1$, $\delta=1$ and $G\in\mathfrak{m}^{3}$.
  2. (2)
    If $k$ is even, then the same argument we used above yields that
    $F_{\alpha}=x_{0}x_{1}+\dots+x_{k}^{2}+\alpha x_{k}x_{k+1}+\dots+\begin{cases}\alpha x_{d-1}^{2}+\alpha\delta x_{d}^{3}+G&\text{if $d$ is odd}\\ \alpha\delta x_{d}^{3}+G&\text{if $d$ is even}\end{cases}$
    where we can again assume $\alpha=1$, $\delta=1$ and $G\in\mathfrak{m}^{3}$. Then transform $F_{1}$ via the change of variables defined by $x_{k}+x_{k+1}\mapsto x_{k+1}$ and $x_{i}\mapsto x_{i}$ for $i\neq k+1$ to obtain
    $F_{1}=x_{0}x_{1}+\dots+x_{k}x_{k+1}+\dots+\begin{cases}x_{d-1}^{2}+x_{d}^{3}+G&\text{if $d$ is odd,}\\ x_{d}^{3}+G&\text{if $d$ is even.}\end{cases}$
Now, if $d\geq 2$, Lemmas 4.2 and 4.3 imply that $F_{\alpha}=x_{0}x_{1}+P_{d-2}+G_{1}$ where $G_{1}\in(x_{2},\dots,x_{d})^{3}$. If $d-2\geq 2$, we can apply the lemmas again to $P_{d-2}+G_{1}$, and continue recursively until we obtain the following:
$F_{1}=x_{0}x_{1}+\dots+\begin{cases}x_{d-1}^{2}+x_{d}^{3}+G&\text{if }d\text{ odd, where }G\in(x_{d-1}^{3},x_{d-1}^{2}x_{d},x_{d-1}x_{d}^{2},x_{d}^{4}),\\ x_{d}^{3}+G&\text{if }d\text{ even, where }G\in(x_{d})^{4}.\end{cases}$
Let’s study both cases separately.
  1. (1)
    Let $d$ be odd. The general case is easily deduced from the case $d=1$, so let $x=x_{0}$ and $y=x_{1}$ for simplicity. Note that $G=Ax^{3}+Bx^{2}y+Cxy^{2}+Dy^{3}$ where $A,B,C,D\in k[[x,y]],$ and $D$ is not a unit, so
    $F_{1}=$ $\ x^{2}+y^{3}+Ax^{3}+Bx^{2}y+Cxy^{2}+Dy^{3}$
    $=$ $\ (1+Ax+By)x^{2}+((1+D)y+Cx)y^{2}.$
    Let $u:=1+D$, which is a unit. Doing the change of variables $x\mapsto x;y\mapsto u^{-1}(Cx+y),$
    $(1+Ax+Bu^{-1}(Cx+y))x^{2}+yu^{-2}(Cx+y)^{2}=vx^{2}+u^{-2}y^{3},$
    where $v=1+(A+Bu^{-1}C)x+(Bu^{-1}+C^{2}u^{-2})y$ is a unit. Now, observe that $F_{1}$ and $v^{-1}F_{1}$ define the same principal ideal, so we can assume that $F_{1}=x^{2}+wx^{3}$, with $w=(vu^{2})^{-1}$. By Lemma 4.1, there exists a unit $t$ such that $t^{3}=w$, so after applying the change of variables $x\mapsto x;y\mapsto t^{-1}y,$ we finally map $F_{1}$ to $x^{2}+y^{3}.$
  2. (2)
    If $d$ is even, we can then write $F_{1}=x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{3}(1+H)$ where $H$ is not a unit. Let $w=1+H$. Since $F_{1}$ and $w^{-1}F_{1}$ generate the same principal ideal, we can assume that $F_{1}=w^{-1}x_{0}x_{1}+\dots+w^{-1}x_{d-2}x_{d-1}+x_{d}^{3}$. Finally, we can apply the change of variables
    $\begin{cases}x_{2m}\mapsto w^{-1}x_{2m},&m<d/2\\ x_{2m+1}\mapsto x_{2m+1},&m<d/2\\ x_{d}\mapsto x_{d},\end{cases}$
    which maps $F_{1}$ to $x_{0}x_{1}+\dots+x_{d-2}x_{d-1}+x_{d}^{3}.$
In particular, it follows from Theorem 3.17 an extension of Theorem 3.1 in [ES05] to characteristic 2:
Corollary 4.6.
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional singular complete local hypersurface ring of characteristic $p=2$, where $k$ is algebraically closed. Then one of the following is true:
  1. (1)
    $R\cong k[[x_{0},\dots,x_{d}]]/(Q_{d}),$ or
  2. (2)
    $e_{HK}(R)\geq e_{HK}(k[[x_{0},\dots,x_{d}]]/(Q_{d}))$,
where $Q_{d}$ is as defined in Definition 2.1.
4.2. Characteristic 3 case.
A slight modification of Enescu and Shimomoto’s proof is enough to extend Theorem 3.4 in [ES05] to characteristic 3.
Theorem 4.7.
Let $(R,\mathfrak{m},k)$ be a $d$-dimensional singular complete local hypersurface ring of characteristic 3. Then, one of the following is true:
  1. (1)
    $R\cong k[[x_{0},...,x_{d}]]/(\sum_{i=0}^{d}x_{i}^{2}),$ or
  2. (2)
    $e_{HK}(R)\geq e_{HK}(k[[x_{0},\dots,x_{d}]])/(x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3}).$
Proof.
Repeating the proof by Enescu and Shimomoto, if we write $R\cong k[[x_{0},\dots,x_{d}]]/(f)$, and $f$ is not in case $(1)$, then we can deform $f$ to $F_{\alpha}=v_{0}x_{0}^{2}+\dots+v_{d-1}x^{2}_{d-1}+v_{d}x_{d}^{3},$ where the $v_{i}$ are units, and the Hilbert-Kunz multiplicity of $F_{\alpha}$ is less than that of $f$. Hence, the element $v_{d}^{-1}F_{\alpha}$ generates the same ideal as $F_{\alpha}.$ We now apply Lemma 3.3 in [ES05] to solve the equations $w_{i}^{2}=v_{d}^{-1}v_{i},$ for $i<d.$ Then, $v_{d}^{-1}F_{\alpha}=x_{0}^{2}+\dots+x_{d-1}^{2}+x_{d}^{3},$ and the result follows. ∎
4.3. Main theorem in characteristic 2 and 3.
Theorem 4.8 (Strong Watanabe–Yoshida conjecture for complete intersections in characteristic 2 and 3).
Let $d\geq 2,$ and $k$ be a field of characteristic $p=2,3$. If $(R,\mathfrak{m},k)$ is a non-regular complete intersection ring of dimension $d$, and $\hat{R}\not\cong R_{p,d}$, then
$e_{\text{HK}}\left(R\right)\geq e_{\text{HK}}\left(S_{p,d}\right)>e_{\text{HK}}\left(R_{p,d}\right)$ (10)
Proof.
The first inequality is proven as in Theorem 2.17 using Theorem 4.5 for characteristic $2$ and Theorem 4.7 in the case of characteristic $3$. The second inequality is the content of Theorems 3.17 and 3.18. ∎
5. Revisiting low dimension and low multiplicity.
As mentioned in the introduction, the Watanabe–Yoshida conjecture has been solved in complete generality only in low-dimensional cases. In particular, the only instances in which the three statements of the conjecture have been solved are dimensions 2 through 4, with characteristic $p>2$. In dimensions 5 through 7, the Watanabe–Yoshida conjecture’s strong form did not always follow from Aberbach and Enescu’s results since some arguments depended on Enescu and Shimomoto’s theorem.
Furthermore, the characteristic 2 case is hardly ever covered in the literature; the two-dimensional case does cover it, but it was not included as part of the general statement of the Watanabe–Yoshida conjecture until [Yos09].
In this section, we will extend these low-dimensional (and low-multiplicity) cases of the conjecture to characteristic 2, and also show that the strong version of the conjecture follows. In other words, the main result of this section is the following:
Theorem 5.1 (Watanabe–Yoshida conjecture).
Let $p>0$ be prime and $2\leq d\leq 6$. Set
$R_{p,d}:=\mathbb{F}_{p}[[x_{0},...,x_{d}]]/(Q_{d}),$
where
$Q_{d}=\begin{cases}x_{0}x_{1}+x_{2}x_{3}+\dots+x_{d-1}x_{d}&\text{if $d$ is odd},\\ x_{0}x_{1}+x_{2}x_{3}+\dots+x_{d-2}x_{d-1}+x_{d}^{2}&\text{if $d$ is even}.\end{cases}$
Let $(R,\mathfrak{m},k)$ be an unmixed non-regular local ring of dimension $d$ and characteristic $p$. Then
  1. (1)
    $e_{\text{HK}}(R)\geq e_{\text{HK}}(R_{p,d})$,
  2. (2)
    if $e_{\text{HK}}\left(R\right)=e_{\text{HK}}\left(R_{p,d}\right)$, then $\hat{R}\cong R_{p,d}\hat{\otimes}_{\mathbb{F}_{p}}k$, i.e. the completion of $R$ is isomorphic to the coordinate ring of a quadric hypersurface.
  3. (3)
    $e_{\text{HK}}(R_{p,d})\geq 1+c_{d}$,
where $c_{d}$ are the coefficients of the Taylor expansion of $\sec x+\tan x$.
We recall the known formulas for $e_{\text{HK}}\left(R_{p,d}\right)$. In characteristic $p\geq 3$, dimensions $4$ through $6$ where computed in [Yos09], and dimension 7 in [CSA24]. The values of $e_{\text{HK}}\left(R_{2,d}\right)$ are from Theorem 3.10.
$d$ 3 4 5 6 7
$e_{\text{HK}}\left(R_{p,d}\right)$, $p\geq 3$ $\frac{4}{3}$ $\frac{29p^{2}+15}{24p^{2}+12}$ $\frac{17p^{2}+12}{15p^{2}+10}$ $\frac{781p^{4}+656p^{2}+315}{720p^{4}+570p^{2}+270}$ $\frac{332p^{4}+304p^{2}+192}{315p^{4}+273p^{2}+168}$
$e_{\text{HK}}\left(R_{2,d}\right)$ $\frac{4}{3}$ $\frac{5}{4}$ $\frac{8}{7}$ $\frac{9}{8}$ $\frac{16}{15}$
Table 3. Hilbert–Kunz multiplicity of the quadric hypersurfaces in low dimensions.
Remark.
Note that $e_{\text{HK}}\left(R_{p,d}\right)$ is a decreasing function on $p>0$ for $4\leq d\leq 7$.
Lemma 5.2 (Proposition 2.13, [WY00]).
If $x$ is an $R$-regular element, then $e_{\text{HK}}\left(R/xR\right)\geq e_{\text{HK}}\left(R\right)$. In particular, for all $p>2$ and $d\geq 1$, we have that $e_{\text{HK}}\left(R_{p,d}\right)\geq e_{\text{HK}}\left(R_{p,d+1}\right)$.
Proof of Theorem 5.1.
We will focus on the first two statements, the strong part of the conjecture, since it has already been proven up to dimension 7 that $e_{\text{HK}}\left(R_{p,d}\right)\geq 1+c_{d}$.
We prove the result in a series of claims, first solving dimensions 3 and 4, then multiplicity 2 to 4, and finally dimensions 5 and 6.
Claim 5.3.
The strong Watanabe–Yoshida conjecture holds for dimension $d=3,4$ and characteristic $p>0$:
  1. (1)
    Let $d=3$ first. Lemma 3.3 in [WY05] implies that if $e>2$, then $e_{\text{HK}}\left(R\right)>\frac{4}{3}=e_{\text{HK}}\left(R_{2,3}\right)$. So we only have to deal with the case when $R$ is a double point, i.e. $e=2$. In that case, there are two options: either $e_{\text{HK}}\left(R\right)=2$, in whose case the result follows, or $e_{\text{HK}}\left(R\right)<2$, in whose case $R$ is a hypersurface by Corollary 1.10, [WY05]. In this last case, the result follows from Theorem 4.8.
  2. (2)
    Now, let us discuss the four-dimensional case. We will carefully adjust the computations in the proof of Theorem 4.3 [WY05] to obtain sharper bounds that work in the characteristic 2 case.
    1. (a)
      Let $e=2$. The same argument applies; if $e_{\text{HK}}\left(R\right)=2$, the result follows since $2>e_{\text{HK}}\left(R_{2,4}\right)$, and if $e_{\text{HK}}\left(R\right)<2$, then $R$ is hypersurface by Corollary 1.10 [WY05] and the strong Watanabe–Yoshida conjecture follows from Theorem 4.8.
    2. (b)
      Let $e=3$. In the computation in Claim 1 in the proof of Theorem 4.3, [WY05], it is shown that $e_{\text{HK}}\left(R\right)\geq 3\left(v_{s}-\frac{2(s-1)^{4}}{4!}\right)$ for $s\geq 1$. For $s=2$, this means $e_{\text{HK}}\left(R\right)\geq\frac{30}{24}=\frac{5}{4}$. To obtain the strict inequality with the value $\frac{5}{4}$, it is enough by setting $s=2.1$, and using the formula
      $v_{s}=\sum_{n=0}^{\lfloor s\rfloor}(-1)^{n}\frac{(s-n)^{d}}{n!(d-n)!}$ (11)
      from Section 3 of [AE12]. This computation yields that $e_{\text{HK}}\left(R\right)\geq 1.3>\frac{5}{4}.$
    3. (c)
      Let $4\leq e\leq 9$. In the computation in Claim 1 in the proof of Theorem 4.3 in [WY01], it is shown that $e_{\text{HK}}\left(R\right)\geq\frac{(13-e)e}{24}$ for $3\leq e\leq 10$, and $\frac{(13-e)e}{24}>\frac{5}{4}$ whenever $4\leq e\leq 9$.
    4. (d)
      Let $10\leq e\leq 30$. In Claim 2 in the proof of the same result, it is shown that $e_{\text{HK}}\left(R\right)\geq\frac{(78-e)e}{384}$, and $\frac{(78-e)e}{384}>\frac{5}{4}$ for every $10\leq e\leq 30$.
    5. (e)
      Let $e\geq 31$. In this case, by the usual inequalities between the usual multiplicity and the Hilbert–Kunz multiplicity, we have that $e_{\text{HK}}\left(R\right)\geq\frac{e}{d!}\geq\frac{31}{24}>\frac{5}{4}.$
Claim 5.4.
The strong Watanabe–Yoshida conjecture holds for dimension $d\geq 3$, characteristic $p>0$, and multiplicity $e(R)=2,3,4$.
First, observe that we can assume the ring is complete, and therefore the homomorphic image of a Cohen–Macaulay ring.
  1. (1)
    If $e(R)=2$, the strong Watanabe–Yoshida conjecture follows from Corollary 1.10 in [WY01] and Theorems 2.17 and 4.8.
  2. (2)
    As for $e(R)=3$, if $e_{\text{HK}}\left(R\right)\geq\frac{3}{2}$, then $e_{\text{HK}}\left(R\right)>e_{\text{HK}}\left(R_{2,d}\right)$ for $d\geq 3$. If $e_{\text{HK}}\left(R\right)<\frac{3}{2}$, then $R$ is Gorenstein by Theorem 2.4 (iii) in [AE12], so if $R$ is not a complete intersection, Theorem 4.1 in [AE12] implies that $e_{\text{HK}}\left(R\right)\geq 13/8>e_{\text{HK}}\left(R_{2,d}\right)$ for all $d\geq 3$ – see Theorem 3.10. Otherwise, $R$ is a complete intersection, and the result follows from Theorem 4.8.
  3. (3)
    Finally, let $e(R)=4$. If $e_{\text{HK}}\left(R\right)\geq\frac{4}{3}$, again $e_{\text{HK}}\left(R\right)>e_{\text{HK}}\left(R_{2,d}\right)$ for $d\geq 4$. Otherwise, we have that $e_{\text{HK}}\left(R\right)<\frac{4}{3}$, and the same argument used in [AE12] implies that $R$ has to be a complete intersection.
Claim 5.5.
The strong Watanabe–Yoshida conjecture holds for dimension $d=5,6$ and characteristic $p>0$.
We revisit the proof of Theorem 5.2 [AE12] — see that paper for details. First, we can assume that the residue field of $R$ is infinite, that $e_{\text{HK}}\left(R\right)>2$, and therefore that $R$ is a domain. We can also assume that it is a domain. Let $J$ be a minimal reduction of $\mathfrak{m}$. In the proof of the theorem, Aberbach and Enescu show that $R$ is either a Cohen–Macaulay ring with minimal multiplicity or $\mu(\mathfrak{m}/J^{*})\leq e-2$, where $J^{*}$ is the tight closure of $J$.
  1. (1)
    In the minimal multiplicity case, either $e_{\text{HK}}\left(R_{p,d}\right)\geq 1.5$ or $R$ is a hypersurface, so in both cases the strong Watanabe–Yoshida conjecture holds in every positive characteristic by Theorems 3.10, 2.17, and 4.8.
  2. (2)
    Otherwise, $\mu(\mathfrak{m}/J^{*})\leq e-2$ and we have to separate between dimension 5 and 6.
    1. (a)
      Let $d=5$. The bounds provided in Aberbach–Enescu are sharp for every multiplicity $e\neq 137$, i.e. $e_{\text{HK}}\left(R\right)>e_{\text{HK}}\left(R_{p,d}\right)$. When $e=137$, the lower bound is $e(R)/d!=\frac{137}{5!}<e_{\text{HK}}\left(R_{p,d}\right)$, so we need to improve the bound in this case, but simply applying Theorem 3.2 in [AE12] and the expression (11) for $r=135$ and $s=1.4$ shows that $e_{\text{HK}}\left(R\right)\geq 4.5$.
    2. (b)
      Let $d=6$. Again, by the inequality between the Hilbert–Kunz multiplicity and the usual multiplicity, we have that for all $e\geq 811$,
      $e_{\text{HK}}\left(R\right)\geq\frac{811}{720}>\frac{810}{720}=\frac{9}{8}.$
      Let us now show the inequality when $5\leq e\leq 811$. For each $s\geq 1$, we define the function
      $G_{s}(e):=e(v_{s}-(e-2)v_{s-1})=-e^{2}v_{s}+(v_{s}+2v_{s-1})e.$
      We know from Theorem 3.2 in [AE12] that $e_{\text{HK}}\left(R\right)\geq G_{s}(e)$ for every $s\geq 1$. We will prove that for every $5\leq e\leq 810$, $G_{s}(e)>\frac{9}{8}$.
      For example, if $s=2.63$ and $5\leq e\leq 9$, it can be easily checked with a computer using (11) that $G_{s}(e)>\frac{9}{8}$. Table 4 shows the value of $s$ that works for each $5\leq e\leq 810$.
      $e$ $s$
      $5\leq e\leq 9$ 2.63
      $10\leq e\leq 16$ 2.31
      $17\leq e\leq 28$ 2.1
      $29\leq e\leq 58$ 1.9
      $59\leq e\leq 296$ 1.6
      $297\leq e\leq 810$ 1.3
      Table 4. Values of $s$ such that $G_{s}(e)>e_{\text{HK}}\left(R_{2,6}\right)$.
Corollary 5.6.
The strong Watanabe–Yoshida conjecture holds for dimension $d\geq 2$ and multiplicity $e(R)\leq 5$.
Proof.
The cases of multiplicity $e(R)=2,3,4$ follow from Claims 5.4 and Watanabe–Yoshida in dimension 2. Thus, let $e(R)=5$. Section 3 of [AE12] shows a characteristic free argument that $e_{\text{HK}}\left(R\right)\geq 1.112$. By Theorem 3.10, $e_{\text{HK}}\left(R_{2,7}\right)=\frac{16}{15}<1.112$. The result follows from the fact that $e_{\text{HK}}\left(R_{p,7}\right)$ is a decreasing function (see Table 3 above), and Lemma 5.2. ∎
Theorem 5.7.
The strong Watanabe–Yoshida conjecture holds for dimension $d=7$ and characteristic $p>2$.
Proof.
If the ring is a complete intersection, this is a consequence of 2.17 and 4.8. If the ring is not a complete intersection, Aberbach and Cox-Steib provide a series of bounds in Section 4 of [CSA24] which are easily seen to be sharp whenever $p>2$ using the fact that $e_{\text{HK}}\left(R_{p,7}\right)$ is a decreasing function on $p>2$ and that $e_{\text{HK}}\left(R_{3,7}\right)\approx 1.056$. ∎
Remark.
Note that $e_{\text{HK}}\left(R_{2,7}\right)\approx 1.066$. All the bounds provided by Aberabach and Cox-Steib in dimension 7 are (positive) characteristic-free, and are enough to cover the characteristic 2 case when $e_{\text{HK}}\left(R\right)\geq 11$, but not when $6\leq e_{\text{HK}}\left(R\right)\leq 10$.